### v1

# High-Capacity and Low-PAPR BICM-OFDM Systems Using Non-Equiprobable and Non-Uniform Constellation Shaping With Clipping and Filtering

Abstract
We address a design of high-capacity and low-peak-to-average power ratio (PAPR) orthogonal frequency-division multiplexing (OFDM) systems based on bit-interleaved coded modulation (BICM) utilizing non-equiprobable and non-uniform (NENU) constellations as well as clipping and filtering (CAF). The proposed constellations are generated using a truncated Gaussian distribution, and the merging of constellation points, where the former creates a non-uniform constellation (NUC), and the latter decreases the number of signal points without compromising the achievable bit-wise mutual information (BMI). Since the proposed constellations are uniquely determined by only the two parameters, each associated with NUC and cardinality, the complexity required for the numerical optimization process can be significantly low. We focus on the constellation design based on one dimension, i.e., pulse amplitude modulation (PAM), which facilitates the reduction of demapping complexity for the BICM receiver. The use of CAF at the transmitter can efficiently reduce the PAPR of OFDM signals; however, it introduces clipping noise that may degrade error rate performance, making the application of clipping noise cancellation (CNC) at the receiver essential. Therefore, we optimize the NENU constellations in the presence of CAF and CNC. Simulation results demonstrate that the combination of constellation shaping with CAF and CNC enables BICM-OFDM systems to simultaneously achieve low PAPR and high spectral efficiency over additive white Gaussian noise (AWGN) as well as frequency-selective Rayleigh fading channels. Furthermore, comparative studies confirm that the proposed system significantly outperforms the single-carrier counterpart (i.e., DFT-precoded BICM-OFDM) in terms of PAPR and bit error rate (BER) performance over fading channels.
I Introduction
To achieve significantly high spectral efficiency under limited spectral resources, the use of quadrature amplitude modulation (QAM) with very large constellation sizes, together with capacity-approaching channel codes, is essential. A common practical approach is to combine bit-interleaved coded modulation (BICM) and the standard QAM, where the QAM is formed by two orthogonal Gray-labeled and equidistant (i.e., uniform) pulse amplitude modulation (PAM) symbols, each corresponding to in-phase (I) and quadrature (Q) channels. In this manner, coding and modulation can be performed in one-dimensional space. Furthermore, BICM allows us to decompose each space into parallel binary channels, where the conventional binary channel codes are applicable. This structure significantly simplifies the complexity of demapping and decoding processes at the receiver [1].
The major drawback of BICM, especially with the standard Gray-labeled PAM symbols, is its achievable information rates that are inferior to the Shannon limit. The standard uniform PAM constellation, with each constellation point spaced equally from its neighbors, differs significantly from Gaussian signals that reach the Shannon limit. This discrepancy leads to a noticeable gap, especially when aiming for a higher target information rate. To reduce this gap, constellation shaping, which essentially controls the distribution of the PAM constellations to Gaussian, plays an important role.
Constellation shaping is classified as geometric and probabilistic shaping. Probabilistic shaping adjusts the probability distribution of the uniform PAM, whereas geometric shaping modifies the placement of constellation points [2, 3]. Probabilistic shaping is typically implemented by introducing a distribution matcher, which entails high computational and storage complexity at the transmitter [4]. Geometric shaping is achieved by altering the locations of constellation points, making it compatible with current systems [5, 6]. The design of a geometrically shaped constellation, often referred to as a nonuniform constellation (NUC), represents a significant area of ongoing research, particularly in conjunction with BICM [7, 8, 9].
In view of practical applications to systems with high spectral efficiency, we focus on the design for NUC of PAM (i.e., one-dimensional constellation) with a large constellation size. It is well known that the two-dimensional constellations can achieve higher performance in terms of the gap from the Shannon limit, but their implementation becomes challenging as the target information rate increases. In fact, existing wireless communication systems that aim for very high spectral efficiency employ modulation schemes based on one-dimensional constellations. For example, the Advanced Television Systems Committee’s third-generation standard (ATSC 3.0) adopts one-dimensional NUC (PAM) for higher-order cases such as $1024$ and $4096$ constellation points [10].
The major drawback of constellation shaping is that it increases peak-to-average power ratio (PAPR) as it attempts to reduce average power for a given maximum amplitude. It causes a significant reduction of power efficiency at the transmitter, as linear amplification of a high PAPR signal is challenging without sacrificing power amplifier efficiency [11]. This high PAPR issue is especially dominant with one-dimensional shaping, as the shape of the resulting two-dimensional constellation is rectangular rather than circular. Therefore, wireless systems that prioritize power amplifier efficiency, such as satellite communications, adopt circular constellations known as amplitude and phase shift keying (APSK), rather than rectangular constellations due to their lower PAPR. Nevertheless, other modern wireless communication systems adopt orthogonal frequency-division multiplexing (OFDM) for its higher spectral efficiency and robustness against fading channels, and since OFDM transmits multiple signals simultaneously, its PAPR turns out to be high regardless of the constellations used for each subcarrier, with or without shaping [12]. In such a framework, it is necessary to apply PAPR reduction techniques developed for OFDM signals, and they should have a high compatibility with constellation shaping.
I-A Our Contributions
In this work, we design an OFDM system with both high spectral efficiency and low PAPR based on the combination of constellation shaping and clipping and filtering (CAF) [13]. Our constellation shaping is based on the use of one-dimensional NUC, where the number of constellation points is reduced from the original PAM consisting of signal points that are powers of two for reduced complexity without noticeable loss of performance. As a result, some constellation points are selected with higher probability than the others, leading to a non-equiprobable constellation. We refer to the resulting constellations as non-equiprobable and non-uniform (NENU) constellations.
Since CAF causes performance loss due to the in-band nonlinear distortion called clipping noise, we also consider the application of clipping noise cancellation (CNC) [14, 15, 16] such that the degradation is significantly mitigated at the receiver side. Overall, the system can achieve both low PAPR and high spectral efficiency, even within the practical framework of OFDM.
We summarize the main contributions of this work as follows:
  • We propose a novel design algorithm for NENU PAM constellations, which combines a geometric design based on a truncated Gaussian distribution with a rate-adaptive constellation merging scheme [^1] In [17], we first presented the concept of merging adjacent signal points for constellations designed according to truncated Gaussian distributions. The algorithm presented in this work is a refined version of [17] in that it can enhance the BMI through the merging process. . Numerical evaluations demonstrate that the proposed scheme achieves performance comparable to the numerically optimized NUC adopted in the ATSC standard.
  • We design BICM-OFDM systems with low PAPR and high spectral efficiency, where low PAPR is achieved by CAF at the transmitter and high spectral efficiency is supported by CNC at the receiver that is compatible with our NENU constellations. We optimize the parameters associated with constellation shaping and PAPR reduction by analyzing bit-wise mutual information (BMI) in the context of CAF/CNC.
  • Through computer simulations, we compare the proposed BICM-OFDM systems with the single-carrier counterpart [18] based on discrete Fourier transform (DFT) precoding that employs the same constellation shaping in terms of the complementary cumulative distribution function (CCDF) of PAPR, demonstrating that the proposed system with proper clipping ratio can achieve lower PAPR than the single-carrier systems in the presence of constellation shaping.
  • Comparison of the resulting bit error rate (BER) performance over additive white Gaussian noise (AWGN) and frequency-selective Rayleigh fading channels shows that the proposed BICM-OFDM system provides noticeable gains in both channels. This is due to the CNC that successfully compensates for the BER degradation caused by CAF with no additional penalty in terms of receiver complexity associated with NENU constellations. It outperforms the single-carrier system with the same constellation over fading channels even with lower PAPR.
I-B Related Work
Several systematic geometric shaping approaches based on Gaussian-like QAM constellations have been proposed in the literature [19, 20, 21] to enhance the achievable information rate while keeping design and implementation complexity modest. Despite their respective design variations, these methods share the common design principle of partitioning a Gaussian distribution into equiprobable sub-regions and assigning a representative constellation point to each sub-region, which was originally suggested in [22]. In our previous work [23], we proposed a systematic approach to designing non-uniform QAM constellations using a truncated Gaussian distribution adjusted by a single parameter. This method flexibly generates the constellation ranging from uniform to Gaussian-like distributions of [22], and we identify the parameter that maximizes the BMI under specified channel signal-to-noise ratios (SNRs). Nevertheless, the designed constellations fail to achieve the shaping gains as high as those of ATSC 3.0 [10], which are optimized based on an exhaustive numerical search. Moreover, the underlying design principles of such optimized NUCs remain largely unexplained, limiting theoretical understanding and systematic reproducibility.
As a low-complexity implementation of probabilistic shaping, a many-to-one mapping approach that does not rely on a distribution matcher can be employed [24, 25, 26]. In this method, redundant bit labels are introduced and they are assigned to a smaller number of constellation points such that the resulting probability mass function (PMF) approximates a Maxwell–Boltzmann (i.e., quantized Gaussian) distribution. Extensions to relatively high-order QAM constellations have recently been investigated in [27, 28, 29]. Since these approaches only require modifications to the bit-to-symbol mapping table, their implementation cost is significantly lower than those based on a distribution matcher. However, they tend to result in lower shaping gains due to their coarse approximation of the target distribution unless a considerably large number of shaping bits are used to adjust the PMF. Recent works [30, 31, 32, 33] have also explored the integration of geometric and probabilistic constellation shaping. Many of these approaches design the constellation geometry and symbol probabilities separately through numerical optimization.
In contrast, our approach offers a unified framework that jointly determines both the positions of constellation points and their probabilities based on a systematic design principle. Furthermore, the application of constellation shaping to OFDM systems with reduced PAPR has not been studied in the literature except for some earlier studies based on trellis shaping [34], which generally requires significant computation for PAPR reduction and causes rate loss for constellation shaping associated with the use of shaping bits.
I-C Organization
The rest of this paper is organized as follows: Section II introduces our NENU constellation construction algorithm, followed by the performance analysis with parameter optimization in Section III. Section IV describes the proposed BICM-OFDM system model employing CAF and CNC, where the achievable information rate in this system model is analyzed. Simulation results are presented in Section V, focusing on the BER performance over practical fading channels and the CCDF of the resulting PAPR. Finally, Section VI concludes this work.
I-D Notation
For a given scalar $A_{k}$, the corresponding bold font $\bf{A}$ represents its row vector form. Sets of real positive numbers, complex numbers, and natural numbers are given by $\mathbb{R}$$_{+}$, $\mathbb{C}$}, and $\mathbb{N}$}, respectively, whereas ${\mathbb{F}}_{2}$ represents the binary field; Calligraphy fonts such as $\mathcal{A}$} represent sets, with their cardinality given by $|{\mathcal{A}}|$; ${\cal N}\left(\mu,\sigma^{2}\right)$ and ${\cal CN}\left(\mu,\sigma^{2}\right)$ represent real-valued Gaussian distribution and circularly symmetric complex Gaussian distribution, respectively, both having mean $\mu$ and variance $\sigma^{2}$. Also, $\mathbb{E}\left[\cdot\right]$ denotes the expectation operation.
II A New Constellation Shaping Algorithm
In this section, we develop a novel systematic design for NENU PAM constellations by incorporating a constellation point merging strategy into the truncated Gaussian-based design. While the capacity of coded modulation (CM) systems increases with the number of distinct constellation points, this is not necessarily the case for BICM, as the optimal constellation cardinality that maximizes the BMI depends on the channel SNR. Therefore, we sequentially merge the adjacent Gaussian sub-regions into larger ones according to the target SNR. In other words, we partition a truncated Gaussian distribution into non-equiprobable sub-regions. We demonstrate that when the two design parameters (i.e., the number of distinct signal points and Gaussian truncation threshold) are jointly optimized, the resulting constellations closely match the numerically optimized NUCs without reducing the number of constellation points adopted in the standard. Unlike the latter, which requires a comprehensive table to specify the constellation points, a pair of parameters uniquely determines our constellation.
The proposed constellation design algorithm consists of two primary steps. We first formulate an initial constellation using a truncated Gaussian distribution and then merge the adjacent points. This algorithm thus facilitates a flexible constellation design in terms of the number of constellation points and their distribution, making it suitable for rate-adaptive designs in BICM systems.
II-A Geometric Design Based on Truncated Gaussian CDF
We first review an $M$-ary PAM constellation design based on the truncated Gaussian distribution [23]. Due to the symmetry of constellation, it is sufficient to consider the constellation points within the positive region. Let $\mathcal{M}_{+}\triangleq\{1,2,\cdots,M/2\}$ denote a set of constellation indices in this region. We define the set of corresponding constellation points as $\mathcal{A}_{+}\triangleq\{a_{k}\in{\mathbb{R}}_{+}\,|\,k\in{\mathcal{M}}_{+}\}$, with the elements ordered in the ascending order, i.e., $0<a_{1}\leq a_{2}\leq\cdots\leq a_{M/2}$. The constellation points are labeled by the binary reflected Gray code (BRGC), where the adjacent points differ only by a single bit. The most significant bit (MSB) is assigned as the sign bit such that $a_{k}$ and $a_{k}$ share the identical bit labels except for the MSB. Without loss of generality, we assume that MSB is 0 for the elements of $\mathcal{A}_{+}$. The resulting constellation including the negative region is given by $\mathcal{A}=\mathcal{A}_{+}\cup\mathcal{A}_{-}$, where $\mathcal{A}_{-}\triangleq\{-a_{k}\,|\,k\in{\mathcal{M}}_{+}\}$.
Since each PAM constellation point is selected with an equal probability, for a given target probability distribution, their locations are determined such that their cumulative distribution function (CDF) values are evenly distributed. Specifically, since the initial target distribution is Gaussian, we consider a random variable $X\sim{\cal N}\left(0,\frac{1}{2}\right)$. The CDF of $X$ is given by
$F_{X}(x)=\frac{1+\operatorname{erf}(x)}{2},$ (1)
where $\operatorname{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt$ is the error function. Since the range of $X$ is unbounded, we truncate its range within [b,b][-b,b] for a given parameter $b>0$. Accordingly, the range of the CDF in the positive region should be modified as $\left[\frac{1}{2},F_{X}(b)\right]$. Thus, the tentative value $\hat{a}_{k}$, which corresponds to the point $a_{k}$ before energy normalization, is determined by the following equality:
$F_{X}(\hat{a}_{k})=\frac{1}{2}+l_{k}\frac{F_{X}(b)-\frac{1}{2}}{M},$ (2)
where
$l_{k}=2k-1,\quad\,\text{ for }\,k\in\mathcal{M}_{+}$ (3)
is the integer value unique to the $k$th constellation point.
Let us define $\rho$ as
$\rho\triangleq\operatorname{erf}(b),\quad\rho\in(0,1].$ (4)
We may then express $\hat{a}_{k}$ in (2) using (1) as
$\hat{a}_{k}=\phi_{M}(l_{k};\rho)\triangleq\operatorname{erf}^{-1}\left(\rho\frac{l_{k}}{M}\right),$ (5)
where $\phi_{M}:\mathbb{N}\to\mathbb{R}_{+}$ is a constellation mapper for a given parameter $\rho$.
Finally, scaling the magnitude of $\hat{a}_{k}$ in accordance with the energy constraint leads to the desired constellation point $a_{k}$:
$a_{k}=\sqrt{\frac{E_{s}}{2}}\frac{\hat{a}_{k}}{\sqrt{\frac{2}{M}\sum_{k=1}^{M/2}\hat{a}^{2}_{k}}},$ (6)
where $E_{s}$ is the energy in terms of QAM symbol.
This design algorithm uniquely determines the resulting constellation using a single parameter $\rho\in(0,1]$ for a given modulation order $M$. When $\rho=1$ (i.e., without truncation), the resulting constellation coincides with the Gaussian-like PAM proposed in [21]. In contrast, we may express the Maclaurin series of $F_{X}(x)$ up to the first order as
$F_{X}(x)\approx\frac{1}{2}+\frac{1}{\sqrt{\pi}}x,$
which indicates that $F_{X}(x)$ linearly increases (i.e., $X$ becomes a uniform distribution) for $X\in[-b,b]$ in the limit of $b\to 0$ ($\rho\to 0$). For this reason, we denote $\rho=0$ as the standard PAM in what follows.
Figure 1: Constellation point selection based on equiprobable partitioning of the Gaussian CDF ($\rho=1$) with $M=32$. Due to the symmetry, only the positive region is shown.
The relationship between the CDF values and the selected constellation points $\{\hat{a}_{k}\}$ in the case of $M=32$ and $\rho=1$ (i.e., without range truncation) is illustrated in Fig. 1. As shown in the figure, the values $\phi_{M}(2(k-1);\rho)$ and $\phi_{M}(2k;\rho)$ serve as the boundary of the $k$th equiprobable sub-region, which are depicted as the corresponding gray dashed lines. Each sub-region is assigned a single bit label, and the representative constellation point $\hat{a}_{k}$ is then selected to match the midpoint CDF value (i.e., $\phi_{M}(l_{k};\rho)$) of its corresponding region.
II-B Constellation Point Merging Process
Since the optimal constellation cardinality that maximizes the BMI varies depending on the channel SNR, we consider dynamically adjusting the number of constellation points by incorporating a constellation point merging strategy into the truncated Gaussian-based design described above.
We define $N$ as the number of distinct points in $M$-ary PAM symbol with $N\leq M$, which means that a PAM constellation with $M$ distinct labels has only $N$ distinct constellation points. We denote this constellation as $(M,N)$-PAM in this work. The construction of $(M,N)$-PAM is given by the following recursive principle: adjacent point pairs in $\hat{\mathcal{A}}_{+}\triangleq\{\hat{a}_{k}\,|\,k\in{\mathcal{M}_{+}}\}$ with the minimum Euclidean distance are merged sequentially until the number of distinct points reaches the specified size $N$. The merging of the two symmetrical pairs in the positive and negative regions occurs simultaneously, leading to a reduction of two constellation points at each merging step. Therefore, for an initial modulation order of $M$, the procedure can be executed a maximum of M/21M/2-1 times in principle, ultimately leading to the standard BPSK constellation (when $N=2$).
Each merging operation combines the corresponding sub-regions into a new, unified sub-region, and the representative constellation point is placed to match the midpoint CDF value of this modified sub-region. Now, suppose that for a given parameter $\rho$, the constellation points $\hat{a}_{i}=\phi_{M}(l_{i};\rho)$ and $\hat{a}_{j}=\phi_{M}(l_{j};\rho)$ (with $i\neq j$) have the minimum Euclidean distance at a certain merging step. Then, these two points are merged into a single point by updating their indices and amplitude values as222 Since the initial indices $\{l_{k}\}$ are set to be positive odd numbers according to (3), the first merging step always applies to the two positive odd numbers. As a result, the merged index $l^{\prime}$ is guaranteed to be a natural number.
$\hat{a}_{i}=\hat{a}_{j}:=\phi_{M}\left(l^{\prime};\rho\right),\quad l^{\prime}:=\frac{l_{i}+l_{j}}{2}.$ (7)
After this process, the multiple bit labels originally associated with $\hat{a}_{i}$ and $\hat{a}_{j}$ are mapped to the identical signal amplitude. The proposed constellation shaping algorithm for a given pair of initial modulation order $M$ and the number of distinct points $N$ after the merging process is summarized in Algorithm 1, where $\mathcal{L}$ corresponds to the set of the constellation indices in the positive region that have unique constellation point locations. The algorithm thus terminates when $|\mathcal{L}|=N/2$.
Figure 2: Constellation point selection based on non-equiprobable partitioning of the Gaussian CDF ($\rho=1$) with $M=32$ and $N=10$.
Algorithm 1 Construction of $(M,N)$-PAM
0: Modulation order $M$; The number of distinct points NN; Shaping parameter $\rho\in(0,1]$
0: Constellation set $\hat{\mathcal{A}}_{+}$
[…]
3:while $M>N$ do
4:  $(i,j)\leftarrow\arg\min_{(i,j)}|\hat{a}_{i}-\hat{a}_{j}|$ for $1\leq i<j\leq M/2$
5:  $\mathcal{L}\leftarrow\mathcal{L}\setminus\{l_{i},l_{j}\}$
6:  $\hat{\mathcal{A}}_{+}\leftarrow\hat{\mathcal{A}}_{+}\setminus\{\hat{a}_{i},\hat{a}_{j}\}$
7:  $l^{\prime}\leftarrow\frac{l_{i}+l_{j}}{2}$
8:  $\hat{a}_{k}\leftarrow\phi_{M}\left(l^{\prime};\rho\right)$
9:  $\mathcal{L}\leftarrow\mathcal{L}\cup\{l^{\prime}\}$
10:  $\hat{\mathcal{A}}_{+}\leftarrow\hat{\mathcal{A}}_{+}\cup\{\hat{a}_{k}\}$
11:  $M\leftarrow M-2$
12:end while
13:return  $\hat{\mathcal{A}}_{+}$
II-C Example
TABLE I: The $l_{k}$ indices of constellation points $\hat{\mathcal{A}}_{+}$ for the $(32,N)$-PAM with $\rho=1$ (no truncation). The second row indicates the bit labels. The initial amplitude values are shown in the third row, while the values after the merging process are shown in brackets. The number in the leftmost column represents that of distinct points $N$.
$k$ 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16
bit label 00000 00001 00011 00010 00110 00111 00101 00100 01100 01101 01111 01110 01010 01011 01001 01000
$\hat{a}_{k}$ .028 .083 .139 .196 .255 .315 .377 .443 .512 .587 .669 .762 .870 1.003 1.185 1.523
32 1 3 5 7 9 11 13 15 17 19 21 23 25 27 29 31
30 2 [.055] 5 7 9 11 13 15 17 19 21 23 25 27 29 31
28 2 6 [.168] 9 11 13 15 17 19 21 23 25 27 29 31
26 2 6 10 [.284] 13 15 17 19 21 23 25 27 29 31
24 2 6 10 14 [.410] 17 19 21 23 25 27 29 31
22 2 6 10 14 18 [.549] 21 23 25 27 29 31
20 2 6 10 14 18 22 [.714] 25 27 29 31
18 4 [.111] 10 14 18 22 25 27 29 31
16 4 12 [.346] 18 22 25 27 29 31
14 4 12 18 22 26 [.932] 29 31
12 4 12 20 [.627] 26 29 31
10 8 [.225] 20 26 29 31
8 8 20 26 30 [1.317]
6 8 20 28 [1.085]
4 8 24 [.813]
2 16 [.477]
The integer representations $\{l_{k}\}$ of the constellation points for the positive region $\hat{\mathcal{A}}_{+}$ obtained by Algorithm 1 are listed in TABLE I in the case of $(32,N)$-PAM starting from $N=32$ to $2$, where the initial constellation points are derived from the original Gaussian distribution (i.e., $\rho=1$). The bit labels are based on BRGC. Therefore, the bit label of the negative region $a_{k}$ is identical to that of $a_{k}$ with MSB replaced by $1$. Focusing on the first merging step from $N=32$ to $30$, the initial constellation points near the origin (e.g., $\hat{a}_{1},\hat{a}_{2},\hat{a}_{3},\hat{a}_{4}$) exhibit very small Euclidean distances. When the adjacent point pair $\hat{a}_{1}$ and $\hat{a}_{2}$ is merged, they are modified into a single signal point but with two distinct bit labels. Then, these merged symbols become indistinguishable from each other, but they become more separable from the remaining symbols (e.g., $\hat{a}_{3},\hat{a}_{4}$). The conventional Gray labeling ensures that the interference from adjacent symbols results in only a single-bit error. However, in high-order constellations under low SNR conditions, the interference from non-adjacent symbols become likely as well, resulting in multiple-bit errors. The proposed constellation merging strategy effectively mitigates this issue by adaptively controlling the balance of density in terms of geometry and probability.
Fig. 2 illustrates the constellation point selection process based on the proposed merging scheme applied to the resulting constellation of Fig. 1, with the number of distinct points set to $N=10$. As shown in the figure, the original equiprobable sub-regions (gray dashed lines) are merged into larger sub-regions, depicted as blue dashed lines. Each representative constellation point is then selected in the same manner as that in Fig. 1, but based on the midpoint CDF value of the modified sub-region. This merging process corresponds to a non-equiprobable partitioning of the reference Gaussian distribution, according to the number of bit labels assigned to each constellation point.
We note that constellation point merging or many-to-one mapping in [26, 28, 29] refers to a low-complexity technique for implementing probabilistic shaping, where redundant bit labels are assigned to control the probability mass function (PMF) of the constellation to achieve shaping gain. In contrast, the proposed approach utilizes constellation point merging as an internal tool of geometric constellation design for the purpose of avoiding excessive density in the constellation. Therefore, this modification does not affect the complexity of the transmitter.
II-D Comparison With NUC of ATSC 3.0
Figure 3: Comparison between the proposed QAM consisting of two (32,20)-PAM constellations (shown as green circles) and the equivalent one-dimensional NUC adopted in ATSC 3.0 (shown as blue diamonds). Both constellations are optimized for BICM with an information rate of 8.333 bits per symbol (code rate of $3/5$).
The proposed QAM constellation points consisting of two $(32,20)$-PAM constellations where the parameter $\rho$ is optimized to maximize the BMI at the information rate corresponding to the code rate of $3/5$ (discussed in the next section) are depicted as green circles in Fig. 3. For comparison, those of the equivalent $1024$-point NUC adopted in ATSC 3.0 are plotted as blue diamonds. While the ATSC NUC individually optimizes all $32$ points per each dimension, many points around the origin are almost overlapping, resembling a smaller number of distinct points. Interestingly, the proposed constellation, systematically derived by Algorithm 1 and governed by only the two design parameters $N$ and $\rho$, exhibits a remarkably similar configuration to the NUC of ATSC 3.0. The performance analysis and optimization of these parameters leading to the above result are presented in the following section.
III Performance Analysis
This section analyzes the performance that can be achieved by the proposed constellation design through parameter optimization, focusing on achievable gain and demapping complexity. We compare the results with other constellations found through numerical optimizations, which require significantly higher complexity than the proposed scheme. From Section II-B, for a given modulation order $M$, the proposed constellation is specified by the two parameters, $N$ and $\rho$. We denote the resulting constellation points located in both positive and negative regions by ${\cal A}\left(M,N,\rho\right)$ in what follows.
III-A Total BMI
To identify the constellations that are expected to perform well, it is necessary to evaluate their achievable information rates. In BICM systems, the achievable information rate is defined by the BMI [1]. For a given $M$-ary PAM constellation $\cal A$}, the total BMI at a specified channel SNR $\gamma$ is given by
$C\left(\mathcal{A},\gamma\right)=n-\sum_{i=0}^{n-1}{\mathbb{E}_{Y,b}\left[\log_{2}\frac{\sum_{x\in\mathcal{A}}p_{Y\mid X}(Y\,|\,x)}{\sum_{x\in\mathcal{A}_{b}^{(i)}}p_{Y\,|\,X}(Y\,|\,x)}\right]},$ (8)
measured in bits per dimension, where $n=\log_{2}M$ is the total number of binary channels and $\mathcal{A}_{b}^{(i)}\subset{\cal A}$ denotes the set of constellation points in the $i$th binary channel with the bit specified by $b\in\mathbb{F}_{2}$. The conditional probability density function (PDF) $p_{Y|X}(y\,|\,x)$ is defined as
$p_{Y|X}(y\,|\,x)=\frac{1}{\sqrt{\pi N_{0}}}e^{-\frac{(y-x)^{2}}{N_{0}}},$ (9)
where $N_{0}$ represents the power spectral density of AWGN, which is related to $\gamma$ by $\gamma=E_{s}/N_{0}$.
III-B Parameter Optimization
For a given pair of modulation order $M$ and operating SNR $\gamma$, the optimization problem in our NENU constellation is formulated as
$\max_{N,\rho}$ $C\Bigl{(}{\cal A}\left(M,N,\rho\right),\gamma\Bigr{)},$ (10)
subject to $N\in\{2,4,\ldots,M\}$
$\rho\in[0,1].$
By evaluating the BMI over all parameter pairs of $(N,\rho)$, the optimal constellation can be identified. (In practice, we may apply the quantized value of $\rho$ with the step size of $0.01$.) Note that when the channel is modeled as AWGN, the expectation over the Gaussian noise in the BMI calculation (8) can be efficiently computed using Gaussian–Hermite quadrature (GHQ) [35].
III-C Loss From the Shannon Limit
Figure 4: Loss in achievable information rate from the Shannon limit for the BICM systems using the proposed constellations with various modulation orders.
The gain achieved by constellation shaping is typically assessed by the difference between the achievable information rate and the Shannon capacity, known as the loss from the Shannon limit [36]. When the two identical $M$-ary PAM constellations form $M^{2}$-ary QAM, the loss from the Shannon limit is defined as a function of SNR $\gamma$ by
$L\left({\cal A},\gamma\right)\triangleq\operatorname{log_{2}}\left(1+\gamma\right)-2\,C\left({\cal A},\gamma\right)$ (11)
in bits per symbol, where the second term corresponds to the total BMI of the two-dimensional QAM constellation.
The loss from the Shannon limit of the proposed shaping scheme is shown in Fig. 4 for various modulation orders ranging from $M^{2}=16$ to $M^{2}=1\,048\,576$. Each plotted point is obtained by selecting the optimal parameter set $\left(N,\rho\right)$ at a given SNR. As a comparison, the minimum loss achieved by the uniform $M^{2}$-QAM with the optimized modulation order is indicated as a gray dashed line. At low SNRs, the performance of the proposed shaping scheme coincides with that of conventional QAM, due to the optimality of low-order constellations such as QPSK and uniform 16-QAM in this SNR regime. However, as the SNR increases, the gap between the BMI of standard QAM and the Shannon limit becomes noticeable. In contrast, the proposed NENU constellation consistently benefits from higher modulation orders across all SNRs. This indicates that, due to its flexible control over the number of signal points, properly designed NENU constellations successfully avoid excessive geometric density that would otherwise degrade the BMI. In fact, the proposed NENU constellation enhances the BMI because of an effect similar to probabilistic shaping.
III-D Comparison With Other Geometric Shaping Schemes
Figure 5: Loss from the Shannon limit for the BICM systems with several different modulation schemes.
We next compare the proposed NENU constellations with other constellations. For this purpose, we fix the modulation order as $M^{2}=1024$ in what follows. We consider the following numerically optimized constellations for comparison:
  • The one-dimensional (1D) $1024$-NUC of ATSC 3.0.
  • The two-dimensional (2D) $1024$-NUC developed in [7].
  • The uniform $M^{2}$-ary QAM with $M^{2}\in\{64,256,1024\}$
  • The one-dimensional shaping without constellation merging, i.e., ${\cal A}\left(M,M,\rho\right)$ with $M^{2}\in\{64,256,1024\}$ (indicated as shaped QAM).
  • The proposed one-dimensional NENU constellation ${\cal A}\left(32,N,\rho\right)$.
The results are compared in Fig. 5. We first observe that the optimal modulation order in BICM is contingent upon the channel SNR. The shaped QAM without constellation merging has been confirmed to demonstrate improved performance compared to uniform QAM [23]. However, due to the constraint of constellation sizes to powers of two, these schemes exhibit significant gaps from the Shannon limit at specific SNR regions, particularly around 15 dB and 20 dB. In the case of the proposed shaping scheme, these gaps can be maintained at a minimal level across a wide range of SNR. This suggests that employing shaping with a flexible constellation size is advantageous for rate-adaptive BICM design. The gain achieved through the proposed shaping scheme is comparable to that of the one-dimensional NUC in ATSC 3.0 across all the SNRs, even though the latter is obtained by a more complex optimization process.
When comparing the results with the two-dimensional constellation of [7], performance gaps of approximately $0.08$ bits per symbol are observed, highlighting the limitations inherent in one-dimensional constellation design. However, one-dimensional modulation schemes offer practical advantages in very high data-rate regimes, since the demapping process for two-dimensional constellations requires a considerable increase in computational cost with higher modulation orders, as is discussed in the next subsection.
III-E Demapping Complexity
Figure 6: The optimal values of constellation size $N$ and design parameter $\rho$ according to the channel SNR ($M=32$).
For efficient decoding, it is necessary to compute bit log-likelihood ratio (LLR) and this dominates the computational complexity at the receiver especially when irregular constellations are adopted. In what follows, we discuss the demapping complexity for general QAM constellations  [^3] There have been several attempts to simplify the soft-output demappers for both one-dimensional and two-dimensional non-uniform constellations [37, 38, 39, 40], but they rely on specific geometric structures of constellations. .
Let us assume that the transmitted symbols are selected from a set of specific complex-valued (i.e., two-dimensional) constellation set $\mathcal{A}$, with the number of points represented by $|\mathcal{A}|$. The LLR value for the $i$th bit level is computed as
$L_{i}$ $=\log\left[\frac{\sum_{x\in\mathcal{A}_{0}^{(i)}}\exp\left(-\frac{|y-x|^{2}}{N_{0}}\right)}{\sum_{x\in\mathcal{A}_{1}^{(i)}}\exp\left(-\frac{|y-x|^{2}}{N_{0}}\right)}\right]$
$\approx\frac{1}{N_{0}}\left(\min_{x\in\mathcal{A}_{1}^{\left(i\right)}}\left\{|y-x|^{2}\right\}-\min_{x\in\mathcal{A}_{0}^{\left(i\right)}}\left\{|y-x|^{2}\right\}\right),$ (12)
where $y\in{\mathbb{C}}$ denotes the received complex-valued symbol. We focus only on the number of required multiplication operations as the complexity associated with comparison and addition operations is considered negligible in comparison to that of multiplication operations.
On the other hand, if $\mathcal{A}$ is a set of one-dimensional constellation points, the LLR value can be expressed as
$L_{i}$ $\approx\frac{1}{N_{0}}\left(\min_{x\in\mathcal{A}_{1}^{\left(i\right)}}\left\{(y-x)^{2}\right\}-\min_{x\in\mathcal{A}_{0}^{\left(i\right)}}\left\{(y-x)^{2}\right\}\right),$ (13)
where $y\in{\mathbb{R}}$ is the received real-valued symbol and $x$ is chosen from one-dimensional (i.e., real-valued) constellation.
In the case of a two-dimensional constellation, from (12) it is necessary to compute the squared Euclidean distances between the received symbol and each of the $|\mathcal{A}|$ constellation points. This computation entails performing $2|\mathcal{A}|$ real multiplications. Furthermore, the cardinality of the two-dimensional constellations is $|\mathcal{A}|=M^{2}$. On the other hand, in the case of a one-dimensional constellation, calculating (13) requires $|\mathcal{A}|$ real multiplications, and the corresponding cardinality is $|\mathcal{A}|=M$. Therefore, as $M$ increases for higher spectral efficiency, the use of a one-dimensional constellation should be critical in view of practical implementation. This reasoning may elucidate the rationale behind the adoption of one-dimensional constellations for high-order scenarios in practical wireless communication systems such as ATSC 3.0, despite its lower achievable gain.
The proposed NENU constellation also offers a more efficient demapping implementation due to its reduced constellation size from $M$, since the demapping complexity is proportional to the actual number of constellation points $N$. With the modulation order set as $M=32$, the optimal constellation size $N$ and the design parameter $\rho$ as a function of channel SNR $\gamma$ are illustrated in Fig. 6. It is observed that the required number of distinct signal points decreases as the SNR decreases. It thus demonstrates that the proposed NENU constellation not only achieves near-optimal gain but also minimizes the demapping complexity.
IV Application to BICM-OFDM With CAF and CNC
The design of wireless signals with low PAPR continues to be a prominent and challenging field of research [41, 42, 43, 44], and a major practical drawback of constellation shaping is its increasing PAPR as it targets Gaussian distribution. On the other hand, BICM-OFDM is a de facto standard for high data rate wireless communications over frequency-selective fading channels, but due to its multicarrier nature, its signal has high PAPR regardless of the constellations employed. Thus, PAPR reduction should be necessary for better amplifier efficiency. Among many PAPR reduction approaches developed for OFDM, we suggest employing CAF [13], as it is directly applicable to the standard OFDM systems. The major drawback is its nonlinear distortion, which should be compensated for at the receiver. Therefore, we also suggest using CNC [14], which allows us to take advantage of one-dimensional constellations for efficient implementation [15]. We demonstrate that the proposed NENU constellations can be flexibly optimized for such systems, achieving notable gains under practical scenarios.
In what follows, for simplicity of notation, we define the set of QAM symbols composed of the two NENU PAM constellations as
${\cal Z}_{M,N,\rho}\triangleq\{X_{I}+jX_{Q}\,|\,X_{I},X_{Q}\in{\cal A}\left(M,N,\rho\right)\}.$
IV-A OFDM Transmitter With CAF
Figure 7: Block diagram of the BICM-OFDM transmitter with CAF. CAF is implemented through a sequence of operations: oversampling using a $JN_{c}$-point IDFT, soft envelope limiter, and out-of-band signal removal.
We follow the description of [13] for CAF and that of [15] for CNC with an extension to coded OFDM. The block diagram of the OFDM transmitter with CAF is illustrated in Fig. 7. The information sequence is first encoded by a binary channel code with code rate $R_{c}\in(0,1]$. The codeword is bit-wise interleaved and then mapped onto the corresponding NENU constellations. Let $\mathbf{Z}=\left(Z_{0},Z_{1},\ldots,Z_{N_{c}-1}\right)\in\mathcal{Z}_{M,N,\rho}^{N_{c}}$ denote a vector of QAM symbols carried by each $N_{c}$-subcarrier OFDM signal, where $Z_{k}=X_{I,k}+jX_{Q,k}$ with $X_{I,k}$ and $X_{Q,k}$ representing NENU PAM symbol on the $k$th subcarrier. The $J$-times oversampled OFDM symbol $\mathbf{S}=(S_{0},S_{1},...,S_{JN_{c}-1})\in{\mathbb{C}}^{JN_{c}}$ is generated by performing $JN_{c}$-point inverse discrete Fourier transform (IDFT) on the zero-padded symbol vector $\mathbf{Z}^{\prime}=(\mathbf{Z},\underbrace{0,0,\ldots,0}_{(J-1)N_{c}})\in{\mathbb{C}}^{JN_{c}}$ as follows:444 In practice, DFT and IDFT are implemented using the fast Fourier transform (FFT) algorithm by setting the number of points $JN_{c}$ as the powers of two.
$S_{n}=\frac{1}{\sqrt{JN_{c}}}\sum_{k=0}^{JN_{c}-1}Z^{\prime}_{k}e^{j2\pi k\frac{n}{JN_{c}}},\quad n=0,1,\ldots,JN_{c}-1.$ (14)
To restrict the peak power, these time domain signal samples are clipped by soft-envelope limiter [13] as
$\tilde{S}_{n}=\begin{cases}S_{n},&\text{ for }\left|S_{n}\right|\leq A_{\text{max}},\\ A_{\text{max}}e^{j\arg S_{n}},&\text{ for }\left|S_{n}\right|>A_{\text{max}},\end{cases}$ (15)
where $A_{\text{max}}$ represents the maximum envelope level. We define the clipping ratio $\gamma_{\text{CR}}$ as
$\gamma_{\text{CR}}\triangleq\frac{A_{\text{max}}}{\sqrt{P_{\text{in}}}},$ (16)
where $P_{\mathrm{in}}\triangleq\mathbb{E}\left[\|\mathbf{Z}\|^{2}\right]/N_{c}$ is the average power of the OFDM signal prior to CAF. We then perform $JN_{c}$-point discrete Fourier transform (DFT) on the clipped signals to transform them back to the frequency-domain symbols represented by $\mathbf{U}^{\prime}=\left(U^{\prime}_{0},U^{\prime}_{1},\ldots,U^{\prime}_{JN_{c}-1}\right)\in{\mathbb{C}}^{JN_{c}}$, where
$U^{\prime}_{k}=\frac{1}{\sqrt{JN_{c}}}\sum_{n=0}^{JN_{c}-1}\tilde{S}_{n}e^{-j2\pi k\frac{n}{JN_{c}}},\,\,k=0,1,\ldots,JN_{c}-1.$ (17)
By removing the out-of-band component from these symbols, we obtain the in-band component ${\mathbf{U}}=\left({U}_{0},{U}_{1},\ldots,{U}_{N_{c}-1}\right)\in{\mathbb{C}}^{N_{c}}$ with ${U}_{k}\equiv{U}^{\prime}_{k}$ as the transmit symbols.
The average power of the transmit OFDM signal after the CAF process is given by $P_{\mathrm{av}}\triangleq\mathbb{E}\left[\|\mathbf{{U}}\|^{2}\right]/N_{c}$, i.e., the average signal power that includes the in-band distortion caused by clipping. For a fair comparison in terms of the transmit signal power and the channel noise variance, we define the SNR of an AWGN channel as
$\mathrm{SNR}\triangleq\frac{P_{\mathrm{av}}}{N_{0}}.$ (18)
The above definition of SNR ensures that the signal part is defined as the actual transmit signal including the in-band distortion (clipping noise).
IV-B OFDM Receiver With CNC
Figure 8: Block diagram of the BICM-OFDM receiver with CNC. CNC iteratively estimates the clipping noise based on the regenerated clipped OFDM signal. Each iteration involves estimating the transmitted symbols and applying the CAF process, including $JN_{c}$-point IDFT and DFT.
At the receiver, we consider performing CNC to mitigate the in-band distortion caused by CAF prior to channel decoding. The block diagram of the OFDM receiver with CNC is illustrated in Fig. 8. In the case of an AWGN channel, the received symbol corresponding to the $k$th subcarrier can be expressed as
$\tilde{Z}_{k}=\alpha Z_{k}+D_{k}+W_{k},$ (19)
where $D_{k}\in{\mathbb{C}}$ denotes the distortion component caused by the CAF process and $W_{k}\in{\mathbb{C}}$ represents the Gaussian noise with $W_{k}\sim\mathcal{CN}\left(0,N_{0}\right)$, both corresponding to the $k$th subcarrier. Assuming that the baseband OFDM signal follows a complex Gaussian distribution, the attenuation factor $\alpha$ due to clipping can be expressed as [13]
$\alpha=1-e^{-\gamma_{\text{CR}}^{2}}+\frac{\sqrt{\pi}\gamma_{\text{CR}}}{2}\operatorname{erfc}(\gamma_{\text{CR}}),$ (20)
where $\operatorname{erfc}(x)=1-\operatorname{erf}(x)$. Provided that the length of cyclic prefix is sufficient compared to that of the channel impulse response, the received symbol over a frequency-selective fading channel can be expressed as
$Y_{k}=H_{k}\left(\alpha Z_{k}+D_{k}\right)+W_{k},$ (21)
where $H_{k}$ is the channel coefficient of the $k$th subcarrier. Assuming that the channel coefficients can be perfectly estimated at the receiver, we may apply the zero-forcing (ZF) equalization as
$\tilde{Y}_{k}\triangleq\frac{Y_{k}}{H_{k}}=\alpha Z_{k}+D_{k}+W_{k}^{\prime},$ (22)
where $W_{k}^{\prime}\triangleq W_{k}/H_{k}\sim\mathcal{CN}\left(0,N_{0}/\left|H_{k}\right|^{2}\right)$.
In the CNC process, we iteratively estimate and remove the distortion component ${\mathbf{D}}=\left({D}_{0},{D}_{1},\ldots,{D}_{N_{c}-1}\right)\in{\mathbb{C}}^{N_{c}}$. The CNC algorithm at the $K$th iteration (starting with $K:=0$) is described as follows:
  1. 1.
    Apply the hard decision based on the constellation set $\mathcal{A}({M,N,\rho})$ to the equalized and normalized received signal $Y_{k}^{(K)}/\alpha$ (with $Y_{k}^{(0)}:=\tilde{Y}_{k}$) to estimate the transmitted symbol $\hat{Z}^{(K)}_{k}={\hat{X}_{I,k}^{(K)}}+j{\hat{X}_{Q,k}^{(K)}}$, which is given by
    $\begin{cases}{\hat{X}}_{I,k}^{(K)}=\arg\min_{{X}\in\mathcal{A}({M,N,\rho})}\left({X}-\Re\left\{{{Y}}_{k}^{(K)}\right\}/\alpha\right)^{2},\\ {\hat{X}}_{Q,k}^{(K)}=\arg\min_{{X}\in\mathcal{A}({M,N,\rho})}\left({X}-\Im\left\{{{Y}}_{k}^{(K)}\right\}/\alpha\right)^{2}.\end{cases}$ (23)
  2. 2.
    Apply the same CAF process at the transmitter to obtain the estimated in-band component, which can be modeled as
    $\hat{{U}}_{k}^{(K)}=\alpha\hat{{Z}}_{k}^{(K)}+\hat{{D}}_{k}^{(K)}.$ (24)
  3. 3.
    Calculate the distortion component by
    $\hat{{D}}_{k}^{(K)}=\hat{{U}}_{k}^{(K)}-\alpha\hat{{Z}}_{k}^{(K)},$ (25)
    and subtract it from the original received symbol as
    ${{Y}}_{k}^{(K+1)}$ $=\tilde{Y}_{k}-\hat{{D}}_{k}^{(K)}$
    $=\alpha{Z}_{k}+\left({D}_{k}-\hat{{D}}_{k}^{(K)}\right)+{W}_{k}.$ (26)
  4. 4.
    Set $K:=K+1$. If $K$ reaches the iteration limit $K_{\text{max}}$, output ${{Y}}_{k}^{(K_{\text{max}})}$ and stop. Otherwise, go to Step 1).
Note that each iteration involves CAF at the receiver side, which requires performing $JN_{c}$-point IDFT, clipping, and $JN_{c}$-point DFT at Step 2).
If the majority of the estimated symbols are correct, the residual distortion component ${D}_{k}-\hat{{D}}_{k}^{(K)}$ should diminish as $K$ increases. This cancellation process may be successful when the channel SNR is sufficiently high. Since our target is a high spectral efficiency regime, this high-SNR assumption naturally applies, and thus CNC is expected to work properly, provided that the clipping ratio $\gamma_{\text{CR}}$ is appropriately selected. Although a lower clipping ratio helps reduce PAPR and improve power amplifier efficiency, it introduces more severe nonlinear distortion, which may prevent CNC from functioning properly.
Note that since our NENU constellations are designed based on a one-dimensional constellation, the complexity associated with the hard decision in (23) is significantly lower than that of the two-dimensional constellation, similar to the argument of demapper complexity. This observation also justifies the use of one-dimensional constellations for our BICM-OFDM systems involving CAF and CNC [^5] It is possible to design CNC that takes into account the channel decoder output as well as re-encoder in its loop, but this approach makes not only the receiver complexity and latency prohibitive, but also the appropriate system design more challenging. .
IV-C Parameter Optimization for OFDM with CAF and CNC
IV-C1 System Setup
In the following numerical results, we consider the specific system where the number of subcarriers is $N_{c}=1024$ and each subcarrier is modulated by $32$-PAM per dimension ($1024$-QAM per subcarrier symbol) with $R_{c}=5/6$. The resulting spectral efficiency is $4.166$ bits per dimension, or $8.333$ bits per complex channel use. Based on the preliminary analysis, the clipping ratio of $\gamma_{\text{CR}}=1.5$ is found to be the minimum acceptable value that does not impose malfunctioning of CNC operation666 Optimal clipping ratio may be mathematically analyzed in terms of signal-to-noise plus distortion power ratio (SNDR) when CAF is applied [45], but since our system makes use of CNC at the receiver, its error rate performance depends on several parameters such as the number of subcarriers and constellations [15]. Therefore, it may be mathematically intractable. . Although some residual clipping noise still remains as a result of CNC, the proposed system can be designed to operate properly with $\gamma_{\text{CR}}=1.5$. Both AWGN and fading channels are considered, where in the case of fading channels we adopt frequency-selective Rayleigh fading channel with an uncorrelated four-path equal-power delay profile as a specific example. The proposed NENU constellation used in each system is optimized based on the corresponding channel condition.
IV-C2 Parameter Optimization
We optimize the system based on (10), where the BMI is computed using (8) by observing the received symbol that serves as an input to the channel decoder. Note that for the systems employing CAF and CNC, the channel becomes a mixture of AWGN and residual clipping noise, which may not be described by a tractable mathematical model. Therefore, we compute the resulting BMI based on simulations, making the optimization process more computationally intensive compared to the AWGN channels without CAF.
We consider the following three cases:
  1. A.
    Without CAF, both AWGN and fading channels
  2. B.
    With CAF and CNC, AWGN channel
  3. C.
    With CAF and CNC, fading channel
For each case, the following optimal parameter sets are found through preliminary simulations:
  1. A.
    $N=32$, $\rho=0.88$  (SNR = $26$ dB)
  2. B.
    $N=24$, $\rho=0.93$  (SNR = $27$ dB)
  3. C.
    $N=28$, $\rho=0.97$  (SNR = $33$ dB)
In what follows, we refer to the NENU constellations designed by the above parameter sets as Modes A, B, and C, respectively.
From the above results, we observe that the optimal number of constellation points depends on the channel conditions, which agrees with the trend discussed in Section III. Without CAF, $N=32$ is selected as optimal due to the high spectral efficiency assumption with $R_{c}=5/6$. In contrast, with CAF and CNC, smaller values of $N$ are preferred to better accommodate the residual nonlinear distortion that reduces the effective SNR.
We note that constellation design based on exhaustive numerical optimization algorithms, such as particle swarm optimization (PSO), does not rely on a predefined set of constellation candidates. Consequently, optimizing the objective functions such as the BMI after CAF and CNC processing becomes computationally demanding. On the other hand, the proposed design requires evaluation over a limited number of constellation candidates determined only by the two parameters. It thus reduces the computational burden significantly and enables optimization in an efficient manner, even under such complicated system constraints.
IV-C3 Total BMI
Figure 9: Comparison of BMI values for 32-PAM constellations over an AWGN channel. The following three systems are evaluated: A1) Without CAF, A2) with CAF only, and A3) with CAF/CNC. (Parameters: $\gamma_{\text{CR}}=1.5$ for CAF and $K_{\text{max}}=10$ for CNC.)
Figure 10: Comparison of BMI values for 32-PAM constellations over a frequency-selective Rayleigh fading channel. The following two systems are evaluated: F1) Without CAF and F2) with CAF and CNC. (Parameters: $\gamma_{\text{CR}}=1.5$ for CAF and $K_{\text{max}}=10$ for CNC.)
We evaluate the total BMI over an AWGN channel for the following three systems:
  • A1:
    BICM-OFDM without CAF (no PAPR reduction)
  • A2:
    BICM-OFDM with CAF but without CNC
  • A3:
    BICM-OFDM with CAF and CNC.
The results are compared in Fig. 9, where the parameters of the NENU constellations are fixed as either Mode A or B over the entire channel SNR. The uniform PAM is also shown as a reference. In A1, Mode A demonstrates superior performance compared to the others at the target information rate. On the other hand, in A2, the achievable rates of all constellations significantly deviate from the target as a result of severe clipping noise. Finally, in A3, the information rate is successfully recovered by CNC, allowing the target information rate to be achieved with a lower PAPR compared to A1. Mode B demonstrates superior performance, yielding a greater gain over Mode A. The results thus demonstrate the effectiveness of the proposed constellation design tailored to the BICM-OFDM system with CAF and CNC.
We next consider the frequency-selective Rayleigh fading channel for the following two systems:
  • F1:
    BICM-OFDM without CAF (no PAPR reduction)
  • F2:
    BICM-OFDM with CAF and CNC.
The BMI values are compared in Fig. 10, where the parameters of the NENU constellations are fixed as either Mode A or C over the entire channel SNR, along with the uniform PAM as a reference. It is observed that for F1, the shaping gains achieved by the proposed constellations decrease compared to the results over the AWGN channel. This may be due to the fact that a single modulation scheme is applied across all subcarriers, while they experience different effective SNRs. Nevertheless, in F2, significant gains by the proposed shaping are still observed when compared at the target information rate. Additionally, Mode C (optimized for this system over fading channel) offers enhanced performance over Mode A.
V Simulation Results
We evaluate the system-level performance of the BICM-OFDM systems employing the proposed NENU constellations in the presence of CAF and CNC. As performance measures, we consider the distribution of PAPR as well as BER. For the binary channel code, we utilize the original rate-$1/3$ turbo code with rate-compatible puncturing [46]. The simulation setup is summarized in TABLE II. We compare the three systems:
  • S1:
    Conventional BICM-OFDM without PAPR reduction.
  • S2:
    BICM-OFDM employing CAF and CNC ($\gamma_{\text{CR}}=1.5$).
  • S3:
    BICM-OFDM employing DFT precoding.
Note that S3 is equivalent to a single-carrier system with rectangular pulse shaping filter and thus its PAPR has direct dependence on the constellation [47].
For each system, we evaluate the uniform PAM, the proposed NENU constellations (parameter sets: Modes A, B, and C), and the one-dimensional NUC of ATSC 3.0. For S3, the conventional minimum mean square error (MMSE) detector is implemented at the receiver for the frequency-selective fading channel.
TABLE II: Simulation setup.
OFDM 1024 subcarriers
Modulation 32-PAM per dimension
Coding Punctured turbo code
Decoding Log-MAP (iteration: 15)
Code rate 5/6
Codeword length 30720
Spectral efficiency $8.333$ bits per subcarrier
PAPR reduction CAF ($\gamma_{\text{CR}}=1.5$, $J=4$) or DFT precoding
PAPR evaluation Four-times oversampling
CNC $K_{\text{max}}=10$
Channel AWGN or Rayleigh fading
Fading channel 4-path equal-power delay profile
V-A CCDF of PAPR
Figure 11: Comparison of the CCDF of PAPR among the three systems: conventional OFDM without CAF, OFDM with CAF, and OFDM with DFT precoding (single-carrier). For each system, the three PAM constellations (uniform, Mode A, and Mode B) are evaluated.
The CCDF curves of the PAPR among the three systems are compared in Fig. 11, where the three PAM constellations (uniform as well as NENU with Modes A and B) are compared.
We observe no difference among the three curves plotted for S1 and S2, suggesting that each subcarrier constellation does not affect PAPR distribution for OFDM. On the other hand, in S3, the selection of constellations significantly affects the dynamic range of the transmit signals, indicating that the use of constellation shaping increases the PAPR as expected [48], even though it is still less than the conventional OFDM (S1). In contrast, from the results of S2, we observe that CAF effectively reduces the peak power of OFDM signals, even in the presence of constellation shaping, achieving a PAPR reduction of $5.0$ dB compared to S1 when measured at a CCDF of $10^{-4}$.
V-B BER Comparison Over AWGN Channel
Figure 12: BER comparison over an AWGN channel for the BICM-OFDM systems without CAF (no PAPR reduction) and with CAF/CNC employing three different PAM constellations.
The BER performance over an AWGN channel for the BICM-OFDM systems with different constellations is compared in Fig. 12, evaluating either without CAF (no PAPR reduction) or with CAF and CNC. As for constellations, Modes A and B are applied to $\bf S1$} and $\bf S2$}, respectively. In addition, the uniform PAM and one-dimensional NUC adopted in ATSC 3.0 are applied to both systems. (Note that the results of S3 are omitted as they exhibit similar performance to S1 over the AWGN channel.) In the case of S1, it can be confirmed that the performance gain achieved by Mode A agrees with that estimated from the BMI values in Fig. 9. The performance is also comparable to that achieved by the one-dimensional NUC of ATSC 3.0.
On the other hand, in the case of S2, Mode B achieves a significant gain, yielding an improvement of approximately $2.35$ dB over the uniform PAM and $0.75$ dB over the NUC when evaluated at a BER of $10^{-5}$. This gain, combined with the performance recovery achieved through CNC, significantly reduces the gap from S1 (without PAPR reduction). It is noteworthy that the performance gap between S1 with the uniform PAM and S2 with Mode B is only $0.50$ dB, while they have a significant difference in terms of PAPR. This suggests that the proposed NENU constellation together with CAF and CNC can significantly reduce PAPR without noticeable SNR penalty over the conventional BICM-OFDM system.
V-C BER Comparison Over Fading Channel
Figure 13: BER comparison over a frequency-selective Rayleigh fading channel for the BICM-OFDM systems without CAF (no PAPR reduction), with CAF/CNC, and with DFT precoding, each employing three different PAM constellations.
Finally, we compare BER performance among the three systems S1, S2, and S3 over a frequency-selective Rayleigh fading channel in Fig. 13. As for constellations, Modes A and C are applied to $\bf S1$} and $\bf S2$}, respectively. In addition, the uniform PAM and one-dimensional NUC adopted in ATSC 3.0 are applied to all three systems for comparison. The performance of S2 is found effective over fading channels, achieving a gain of approximately $2.50$ dB with the proposed NENU constellation. On the other hand, the performance of S3 (employing an MMSE detector) is considerably inferior to that of the other OFDM-based systems. It is known that single-carrier signals experience BER degradation when suboptimal detectors, such as MMSE, are employed at the receiver [49, 50]. The results thus indicate that the performance loss associated with DFT precoding is more substantial than that caused by CAF with CNC. Consequently, for PAPR reduction in OFDM systems utilizing constellation shaping, CAF combined with CNC is superior to DFT precoding in terms of both the PAPR reduction effect and BER performance over fading channels.
VI Conclusion
In this work, we developed high-capacity and low-PAPR OFDM-BICM systems based on one-dimensional NENU constellations for shaping gain along with CAF and CNC for efficient signal peak power reduction without noticeable performance degradation. As a result, the developed systems enjoy significantly lower PAPR while achieving comparable performance with the standard PAM-based BICM-OFDM system. The simulation results have demonstrated that significant performance gain can be achieved over the conventional BICM-OFDM with DFT precoding in terms of both PAPR distribution and BER. In our simulations, we have restricted our attention to the design of $32$-PAM (i.e., $1024$-QAM), but its extension to even higher-order constellations targeting even higher information rates is straightforward, considering the simplicity of the optimization process based on only the two distinct parameters.

### v2

@@
 We address a design of high-capacity and low-peak-to-average power ratio (PAPR) orthogonal frequency-division multiplexing (OFDM) systems based on bit-interleaved coded modulation (BICM) utilizing non-equiprobable and non-uniform (NENU) constellations as well as clipping and filtering (CAF).
-The proposed constellations are generated using a truncated Gaussian distribution, and the merging of constellation points, where the former creates a non-uniform constellation (NUC), and the latter decreases the number of signal points without compromising the achievable bit-wise mutual information (BMI).
-Since the proposed constellations are uniquely determined by only the two parameters, each associated with NUC and cardinality, the complexity required for the numerical optimization process can be significantly low.
+The proposed constellations are generated using a truncated Gaussian distribution and the merging of constellation points, where the former creates a non-uniform constellation (NUC), and the latter adjusts the number of signal points for further improving the total bit-wise mutual information (BMI).
+Unlike other exhaustive search-based approaches, the proposed constellations are uniquely determined by only two parameters associated with NUC and cardinality.
+Due to this property of limited degrees of freedom, the complexity required for the numerical optimization process can be significantly low.
 We focus on the constellation design based on one dimension, i.e., pulse amplitude modulation (PAM), which facilitates the reduction of demapping complexity for the BICM receiver.
@@
 Therefore, we optimize the NENU constellations in the presence of CAF and CNC.
-Simulation results demonstrate that the combination of constellation shaping with CAF and CNC enables BICM-OFDM systems to simultaneously achieve low PAPR and high spectral efficiency over additive white Gaussian noise (AWGN) as well as frequency-selective Rayleigh fading channels.
+Simulation results demonstrate that the combination of constellation shaping with CAF and CNC enables BICM-OFDM systems to simultaneously achieve low PAPR and high spectral efficiency over additive white Gaussian noise (AWGN) as well as frequency-selective fading channels.
 Furthermore, comparative studies confirm that the proposed system significantly outperforms the single-carrier counterpart (i.e., DFT-precoded BICM-OFDM) in terms of PAPR and bit error rate (BER) performance over fading channels.
@@
 To achieve significantly high spectral efficiency under limited spectral resources, the use of quadrature amplitude modulation (QAM) with very large constellation sizes, together with capacity-approaching channel codes, is essential.
-A common practical approach is to combine bit-interleaved coded modulation (BICM) and the standard QAM, where the QAM is formed by two orthogonal Gray-labeled and equidistant (i.e., uniform) pulse amplitude modulation (PAM) symbols, each corresponding to in-phase (I) and quadrature (Q) channels.
-In this manner, coding and modulation can be performed in one-dimensional space.
-Furthermore, BICM allows us to decompose each space into parallel binary channels, where the conventional binary channel codes are applicable.
-This structure significantly simplifies the complexity of demapping and decoding processes at the receiver [1].
-The major drawback of BICM, especially with the standard Gray-labeled PAM symbols, is its achievable information rates that are inferior to the Shannon limit.
-The standard uniform PAM constellation, with each constellation point spaced equally from its neighbors, differs significantly from Gaussian signals that reach the Shannon limit.
-This discrepancy leads to a noticeable gap, especially when aiming for a higher target information rate.
-To reduce this gap, constellation shaping, which essentially controls the distribution of the PAM constellations to Gaussian, plays an important role.
-Constellation shaping is classified as geometric and probabilistic shaping.
-Probabilistic shaping adjusts the probability distribution of the uniform PAM, whereas geometric shaping modifies the placement of constellation points [2, 3].
-Probabilistic shaping is typically implemented by introducing a distribution matcher, which entails high computational and storage complexity at the transmitter [4].
-Geometric shaping is achieved by altering the locations of constellation points, making it compatible with current systems [5, 6].
-The design of a geometrically shaped constellation, often referred to as a nonuniform constellation (NUC), represents a significant area of ongoing research, particularly in conjunction with BICM [7, 8, 9].
-In view of practical applications to systems with high spectral efficiency, we focus on the design for NUC of PAM (i.e., one-dimensional constellation) with a large constellation size.
-It is well known that the two-dimensional constellations can achieve higher performance in terms of the gap from the Shannon limit, but their implementation becomes challenging as the target information rate increases.
-In fact, existing wireless communication systems that aim for very high spectral efficiency employ modulation schemes based on one-dimensional constellations.
-For example, the Advanced Television Systems Committee’s third-generation standard (ATSC 3.0) adopts one-dimensional NUC (PAM) for higher-order cases such as $1024$ and $4096$ constellation points [10].
-The major drawback of constellation shaping is that it increases peak-to-average power ratio (PAPR) as it attempts to reduce average power for a given maximum amplitude.
-It causes a significant reduction of power efficiency at the transmitter, as linear amplification of a high PAPR signal is challenging without sacrificing power amplifier efficiency [11].
-This high PAPR issue is especially dominant with one-dimensional shaping, as the shape of the resulting two-dimensional constellation is rectangular rather than circular.
-Therefore, wireless systems that prioritize power amplifier efficiency, such as satellite communications, adopt circular constellations known as amplitude and phase shift keying (APSK), rather than rectangular constellations due to their lower PAPR.
-Nevertheless, other modern wireless communication systems adopt orthogonal frequency-division multiplexing (OFDM) for its higher spectral efficiency and robustness against fading channels, and since OFDM transmits multiple signals simultaneously, its PAPR turns out to be high regardless of the constellations used for each subcarrier, with or without shaping [12].
-In such a framework, it is necessary to apply PAPR reduction techniques developed for OFDM signals, and they should have a high compatibility with constellation shaping.
+A common practical approach is to employ bit-interleaved coded modulation (BICM), which separates modulation and coding by decomposing the transmitted constellation into parallel binary-input channels, to which conventional binary channel codes can readily be applied.
+When combined with standard QAM, the constellation can be decomposed into two Gray-labeled and equidistant (i.e., uniform) pulse amplitude modulation (PAM) constellations, each corresponding to the in-phase (I) and quadrature (Q) components.
+This structure significantly simplifies the complexity of demapping and decoding processes at the receiver [szczecinski2015bit].
+A key challenge in BICM, especially with standard PAM, is its achievable information rates that are inferior to the Shannon limit.
+The standard PAM constellation, with each constellation point spaced equally from its neighbors, differs significantly from Gaussian signals that achieve the Shannon limit.
+This discrepancy leads to a noticeable gap, especially when aiming for a high target spectral efficiency.
+To reduce this gap, constellation shaping, which essentially tailors the distribution of PAM constellations to approximate a Gaussian distribution, plays an important role.
+Constellation shaping is categorized as geometric and probabilistic shaping.
+Probabilistic shaping adjusts the probability distribution of the uniform PAM, whereas geometric shaping modifies the placement of constellation points [steiner2017comparison, qu2019probabilistic].
+Probabilistic shaping is typically implemented by introducing a distribution matcher, which entails high computational and storage complexity at the transmitter [schulte2015constant].
+In contrast, geometric shaping is simply achieved by altering the locations of constellation points, making it compatible with current systems [fay2016overview, singya2021survey].
+Accordingly, the design of a geometrically shaped constellation, often referred to as a non-uniform constellation (NUC), represents a significant area of ongoing research, particularly in conjunction with BICM [sillekens2022high, gumucs2023simplified, chen2022geometrically].
+In view of practical applications to systems with high spectral efficiency, we focus on the design of NUCs based on PAM (i.e., one-dimensional NUCs) with large constellation sizes.
+It is well known that two-dimensional constellations can achieve better performance in terms of the gap from the Shannon limit.
+However, their implementation becomes increasingly challenging as the target information rate increases due to the prohibitively high demapping complexity associated with large constellation sizes.
+In fact, existing wireless communication systems targeting very high spectral efficiency often employ modulation schemes based on one-dimensional constellations.
+For example, the Advanced Television Systems Committee’s third-generation standard (ATSC 3.0) adopts only one-dimensional NUCs for higher modulation orders such as $1024$ and $4096$ [atsc30].
+A major drawback of constellation shaping is the increase in peak-to-average power ratio (PAPR), as it aims to reduce the average power for a given maximum amplitude.
+This leads to a significant reduction of power efficiency at the transmitter, as linear amplification of high PAPR signals is challenging without sacrificing power amplifier efficiency [ochiai2013analysis].
+Nevertheless, the majority of modern wireless communication systems adopt orthogonal frequency-division multiplexing (OFDM) because of its advantages, such as high spectral efficiency and robustness against fading channels.
+Since OFDM transmits multiple signals simultaneously, its PAPR becomes high irrespective of the constellations used for each subcarrier, with or without shaping [ochiai_TCOM_2001].
+In such systems, it is necessary to apply PAPR reduction techniques originally developed for OFDM while ensuring high compatibility with constellation shaping.
 I-A Our Contributions
-In this work, we design an OFDM system with both high spectral efficiency and low PAPR based on the combination of constellation shaping and clipping and filtering (CAF) [13].
-Our constellation shaping is based on the use of one-dimensional NUC, where the number of constellation points is reduced from the original PAM consisting of signal points that are powers of two for reduced complexity without noticeable loss of performance.
+In this work, our target is to implement a BICM-OFDM system that achieves both high spectral efficiency and low PAPR without imposing high complexity at the transmitter side.
+In our shaping approach, a Gray-labeled non-uniform PAM constellation, whose cardinality is a power of two, is first systematically constructed based on a truncated Gaussian distribution characterized by a single parameter.
+The number of constellation points is then sequentially reduced by merging the closest points.
+This merging process improves the achievable information rate at a given signal-to-noise ratio (SNR), while reducing the demapping complexity.
 As a result, some constellation points are selected with higher probability than the others, leading to a non-equiprobable constellation.
 We refer to the resulting constellations as non-equiprobable and non-uniform (NENU) constellations.
-Since CAF causes performance loss due to the in-band nonlinear distortion called clipping noise, we also consider the application of clipping noise cancellation (CNC) [14, 15, 16] such that the degradation is significantly mitigated at the receiver side.
-Overall, the system can achieve both low PAPR and high spectral efficiency, even within the practical framework of OFDM.
+For achieving low PAPR, we adopt clipping and filtering (CAF) [ochiai2002performance] among various PAPR reduction techniques proposed for OFDM.
+While many PAPR reduction schemes impose demanding complexity at the transmitter side, CAF can be implemented in a straightforward manner.
+Furthermore, it is directly applicable to BICM-OFDM systems with any constellation shaping.
+Therefore, combining constellation shaping with CAF is a natural consequence for BICM-OFDM.
+However, CAF inevitably causes performance degradation due to the in-band nonlinear distortion called clipping noise.
+Therefore, we also consider the application of clipping noise cancellation (CNC) [chen2003iterative, sun2021, wachowiak2023clipping] to mitigate this impairment at the receiver.
+To tailor the constellation design to OFDM transmission under CAF and CNC, the two shaping parameters (i.e., the truncation level associated with the initial Gaussian distribution and the number of distinct signal points) are jointly optimized by taking into account the effects of both additive white Gaussian noise (AWGN) and residual clipping noise.
+Overall, the proposed system can simultaneously achieve low PAPR and high spectral efficiency, even within a practical framework of OFDM over frequency-selective fading channels.
 We summarize the main contributions of this work as follows:
-  • We propose a novel design algorithm for NENU PAM constellations, which combines a geometric design based on a truncated Gaussian distribution with a rate-adaptive constellation merging scheme [^1] In [17], we first presented the concept of merging adjacent signal points for constellations designed according to truncated Gaussian distributions.
-The algorithm presented in this work is a refined version of [17] in that it can enhance the BMI through the merging process.
+  • We propose a novel design algorithm for NENU PAM constellations, which combines a geometric design based on a truncated Gaussian distribution with a rate-adaptive constellation merging scheme [^1] In [kurihara2025systematic], we first presented the concept of merging adjacent signal points for constellations designed according to truncated Gaussian distributions.
+The algorithm presented in this work is a refined version of [kurihara2025systematic] in that it can enhance the BMI through the merging process.
 .
 Numerical evaluations demonstrate that the proposed scheme achieves performance comparable to the numerically optimized NUC adopted in the ATSC standard.
-  • We design BICM-OFDM systems with low PAPR and high spectral efficiency, where low PAPR is achieved by CAF at the transmitter and high spectral efficiency is supported by CNC at the receiver that is compatible with our NENU constellations.
-We optimize the parameters associated with constellation shaping and PAPR reduction by analyzing bit-wise mutual information (BMI) in the context of CAF/CNC.
-  • Through computer simulations, we compare the proposed BICM-OFDM systems with the single-carrier counterpart [18] based on discrete Fourier transform (DFT) precoding that employs the same constellation shaping in terms of the complementary cumulative distribution function (CCDF) of PAPR, demonstrating that the proposed system with proper clipping ratio can achieve lower PAPR than the single-carrier systems in the presence of constellation shaping.
-  • Comparison of the resulting bit error rate (BER) performance over additive white Gaussian noise (AWGN) and frequency-selective Rayleigh fading channels shows that the proposed BICM-OFDM system provides noticeable gains in both channels.
-This is due to the CNC that successfully compensates for the BER degradation caused by CAF with no additional penalty in terms of receiver complexity associated with NENU constellations.
+  • We design BICM-OFDM systems with low PAPR and high spectral efficiency, where low PAPR is achieved by CAF at the transmitter and high spectral efficiency is supported by CNC at the receiver that is compatible with our proposed shaping scheme.
+Furthermore, the NENU constellations are optimized in the presence of CAF and CNC.
+  • Through computer simulations, we compare the proposed BICM-OFDM systems with their single-carrier counterparts [myung2006single] based on discrete Fourier transform (DFT) precoding that employs the same constellation shaping in terms of the complementary cumulative distribution function (CCDF) of PAPR, demonstrating that the proposed system with the proper clipping ratio can achieve lower PAPR than the single-carrier systems.
+  • Comparison of the resulting bit error rate (BER) performance over AWGN and frequency-selective fading channels shows that the proposed BICM-OFDM system provides noticeable gains in both channels.
+This is due to the CNC that successfully compensates for the BER degradation caused by CAF with no additional complexity associated with the use of constellation shaping.
 It outperforms the single-carrier system with the same constellation over fading channels even with lower PAPR.
 I-B Related Work
-Several systematic geometric shaping approaches based on Gaussian-like QAM constellations have been proposed in the literature [19, 20, 21] to enhance the achievable information rate while keeping design and implementation complexity modest.
-Despite their respective design variations, these methods share the common design principle of partitioning a Gaussian distribution into equiprobable sub-regions and assigning a representative constellation point to each sub-region, which was originally suggested in [22].
-In our previous work [23], we proposed a systematic approach to designing non-uniform QAM constellations using a truncated Gaussian distribution adjusted by a single parameter.
-This method flexibly generates the constellation ranging from uniform to Gaussian-like distributions of [22], and we identify the parameter that maximizes the BMI under specified channel signal-to-noise ratios (SNRs).
-Nevertheless, the designed constellations fail to achieve the shaping gains as high as those of ATSC 3.0 [10], which are optimized based on an exhaustive numerical search.
+Several systematic geometric shaping approaches based on Gaussian-like PAM constellations have been proposed in the literature [koike2017bit, wiens2020constellation, boutros2018geometric] to enhance the achievable information rate while keeping design and implementation complexity modest.
+Despite their respective design variations, these methods share the common design principle of partitioning a Gaussian distribution into equiprobable subregions and assigning a representative constellation point to each subregion, as originally suggested in [sun1993approaching].
+In our previous work [kurihara2024design], we proposed a systematic approach to designing non-uniform PAM constellations using a truncated Gaussian distribution adjusted by a single parameter.
+This method generates the PAM constellation, whose distribution can be adjusted ranging from Gaussian-like to uniform, and we identify the parameter that maximizes the achievable information rate under specified channel SNRs.
+Nevertheless, the designed constellations fail to achieve shaping gains as high as those of ATSC 3.0 [atsc30], which rely on the optimization through an exhaustive numerical search.
 Moreover, the underlying design principles of such optimized NUCs remain largely unexplained, limiting theoretical understanding and systematic reproducibility.
-As a low-complexity implementation of probabilistic shaping, a many-to-one mapping approach that does not rely on a distribution matcher can be employed [24, 25, 26].
-In this method, redundant bit labels are introduced and they are assigned to a smaller number of constellation points such that the resulting probability mass function (PMF) approximates a Maxwell–Boltzmann (i.e., quantized Gaussian) distribution.
-Extensions to relatively high-order QAM constellations have recently been investigated in [27, 28, 29].
+As a low-complexity implementation of probabilistic shaping, a many-to-one mapping approach that does not rely on a distribution matcher can be employed [yankov2014rate, chang2013symbol, zhou2020enhancement].
+This approach introduces redundant bit labels, which are assigned to a smaller number of constellation points such that the resulting probability mass function (PMF) approximates a Maxwell–Boltzmann (i.e., quantized Gaussian) distribution.
+Extensions to relatively high-order QAM constellations have recently been investigated in [yankov2016constellation, jia2023many, cao2020probabilistic].
 Since these approaches only require modifications to the bit-to-symbol mapping table, their implementation cost is significantly lower than those based on a distribution matcher.
 However, they tend to result in lower shaping gains due to their coarse approximation of the target distribution unless a considerably large number of shaping bits are used to adjust the PMF.
-Recent works [30, 31, 32, 33] have also explored the integration of geometric and probabilistic constellation shaping.
+Recent works [li2024model, soleimanzade2023hybrid, stark2019joint, maneekut2020hybrid] have also explored the integration of geometric and probabilistic constellation shaping.
 Many of these approaches design the constellation geometry and symbol probabilities separately through numerical optimization.
 In contrast, our approach offers a unified framework that jointly determines both the positions of constellation points and their probabilities based on a systematic design principle.
-Furthermore, the application of constellation shaping to OFDM systems with reduced PAPR has not been studied in the literature except for some earlier studies based on trellis shaping [34], which generally requires significant computation for PAPR reduction and causes rate loss for constellation shaping associated with the use of shaping bits.
+We also address the optimization of constellations in the presence of CAF and CNC, which involve complex nonlinear signal processing operations at both the transmitter and receiver.
+Most existing constellation shaping approaches through numerical optimization deal with analytically tractable channels such as AWGN.
+On the other hand, since the proposed approach relies on a predefined family of constellations characterized by limited degrees of freedom, it facilitates efficient optimization even for such channels that cannot be described by explicit mathematical models.
+Lastly, the application of geometric shaping to OFDM systems in conjunction with PAPR reduction techniques has not been well investigated in the literature.
+[^2] Trellis shaping [forney92] can be applied to PAPR and average power reduction of BICM-OFDM systems [ochiai_comm04, yoshizawa2015].
+However, the transmitter complexity required for effective PAPR reduction is considerably higher than CAF, especially when OFDM has a large number of subcarriers.
 I-C Organization
@@
 Section IV describes the proposed BICM-OFDM system model employing CAF and CNC, where the achievable information rate in this system model is analyzed.
-Simulation results are presented in Section V, focusing on the BER performance over practical fading channels and the CCDF of the resulting PAPR.
+Simulation results are presented in Section V, focusing on the BER performance over AWGN and practical fading channels as well as the CCDF of the resulting PAPR.
 Finally, Section VI concludes this work.
 I-D Notation
-For a given scalar $A_{k}$, the corresponding bold font $\bf{A}$ represents its row vector form.
-Sets of real positive numbers, complex numbers, and natural numbers are given by $\mathbb{R}$$_{+}$, $\mathbb{C}$}, and $\mathbb{N}$}, respectively, whereas ${\mathbb{F}}_{2}$ represents the binary field; Calligraphy fonts such as $\mathcal{A}$} represent sets, with their cardinality given by $|{\mathcal{A}}|$; ${\cal N}\left(\mu,\sigma^{2}\right)$ and ${\cal CN}\left(\mu,\sigma^{2}\right)$ represent real-valued Gaussian distribution and circularly symmetric complex Gaussian distribution, respectively, both having mean $\mu$ and variance $\sigma^{2}$.
-Also, $\mathbb{E}\left[\cdot\right]$ denotes the expectation operation.
+For a given scalar $A_{k}$, the corresponding bold font $\mathbf{A}$ denotes its row-vector form.
+Sets of real positive numbers, complex numbers, and natural numbers are given by $\mathbb{R}$$_{+}$, $\mathbb{C}$}, and $\mathbb{N}$}, respectively, whereas ${\mathbb{F}}_{2}$ represents the binary field.
+Calligraphic fonts such as $\mathcal{A}$} denote sets, with their cardinality given by $|{\mathcal{A}}|$, whereas ${\cal N}\left(\mu,\sigma^{2}\right)$ and ${\cal CN}\left(\mu,\sigma^{2}\right)$ represent the real-valued Gaussian distribution and the circularly symmetric complex Gaussian distribution, respectively, both having mean $\mu$ and variance $\sigma^{2}$.
+The expectation operator is represented by $\mathbb{E}\left[\cdot\right]$.
 II A New Constellation Shaping Algorithm
-In this section, we develop a novel systematic design for NENU PAM constellations by incorporating a constellation point merging strategy into the truncated Gaussian-based design.
-While the capacity of coded modulation (CM) systems increases with the number of distinct constellation points, this is not necessarily the case for BICM, as the optimal constellation cardinality that maximizes the BMI depends on the channel SNR.
-Therefore, we sequentially merge the adjacent Gaussian sub-regions into larger ones according to the target SNR.
-In other words, we partition a truncated Gaussian distribution into non-equiprobable sub-regions.
-We demonstrate that when the two design parameters (i.e., the number of distinct signal points and Gaussian truncation threshold) are jointly optimized, the resulting constellations closely match the numerically optimized NUCs without reducing the number of constellation points adopted in the standard.
-Unlike the latter, which requires a comprehensive table to specify the constellation points, a pair of parameters uniquely determines our constellation.
-The proposed constellation design algorithm consists of two primary steps.
-We first formulate an initial constellation using a truncated Gaussian distribution and then merge the adjacent points.
-This algorithm thus facilitates a flexible constellation design in terms of the number of constellation points and their distribution, making it suitable for rate-adaptive designs in BICM systems.
+The proposed design algorithm for NENU constellations consists of two primary steps.
+First, an initial Gray-labeled PAM constellation is derived by partitioning a truncated Gaussian distribution into equiprobable subregions, where each constellation point is assigned according to the midpoint cumulative distribution function (CDF) value of the corresponding region.
+The resulting PAM distribution is controlled by a single parameter that determines the truncation level of the reference Gaussian distribution.
+Next, adjacent subregions are sequentially merged to adjust the number of distinct signal points.
+In that case, the reference Gaussian distribution is partitioned such that the probability associated with a given subregion is proportional to the number of assigned binary labels.
+This algorithm facilitates a flexible constellation design in terms of both the number of constellation points and their distribution, making it suitable for rate-adaptive BICM systems.
+We demonstrate that when the two design parameters (i.e., the number of distinct signal points and the Gaussian truncation level) are jointly optimized, the resulting constellations closely match the optimized NUCs adopted in the ATSC standard.
+Unlike the latter, which requires a comprehensive table to specify the constellation points, only a pair of parameters uniquely determines our constellation.
+As described in Section IV, this “limited degrees of freedom” description turns out to be particularly useful when designing constellations over highly complex channels that may only be modeled by simulation-based approaches.
 II-A Geometric Design Based on Truncated Gaussian CDF
-We first review an $M$-ary PAM constellation design based on the truncated Gaussian distribution [23].
-Due to the symmetry of constellation, it is sufficient to consider the constellation points within the positive region.
+We first review an $M$-ary PAM constellation design based on the truncated Gaussian distribution [kurihara2024design], where each PAM symbol has a binary label ${\bf b}\in{\mathbb{F}}_{2}^{m}$ with $m=\log_{2}M$.
+Due to the symmetry of the constellation, it is sufficient to consider the constellation points within the positive region.
 Let $\mathcal{M}_{+}\triangleq\{1,2,\cdots,M/2\}$ denote a set of constellation indices in this region.
 We define the set of corresponding constellation points as $\mathcal{A}_{+}\triangleq\{a_{k}\in{\mathbb{R}}_{+}\,|\,k\in{\mathcal{M}}_{+}\}$, with the elements ordered in the ascending order, i.e., $0<a_{1}\leq a_{2}\leq\cdots\leq a_{M/2}$.
-The constellation points are labeled by the binary reflected Gray code (BRGC), where the adjacent points differ only by a single bit.
+The constellation points are labeled by the binary reflected Gray code (BRGC) of $m$ bits, where the adjacent points differ only by a single bit.
 The most significant bit (MSB) is assigned as the sign bit such that $a_{k}$ and $a_{k}$ share the identical bit labels except for the MSB.
 Without loss of generality, we assume that MSB is 0 for the elements of $\mathcal{A}_{+}$.
-The resulting constellation including the negative region is given by $\mathcal{A}=\mathcal{A}_{+}\cup\mathcal{A}_{-}$, where $\mathcal{A}_{-}\triangleq\{-a_{k}\,|\,k\in{\mathcal{M}}_{+}\}$.
-Since each PAM constellation point is selected with an equal probability, for a given target probability distribution, their locations are determined such that their cumulative distribution function (CDF) values are evenly distributed.
+The resulting constellation, including the negative region, is given by $\mathcal{A}=\mathcal{A}_{+}\cup\mathcal{A}_{-}$, where $\mathcal{A}_{-}\triangleq\{-a_{k}\,|\,k\in{\mathcal{M}}_{+}\}$.
+Since each PAM constellation point is selected with an equal probability, for a given target probability distribution, their locations are determined such that their CDF values are evenly distributed.
 Specifically, since the initial target distribution is Gaussian, we consider a random variable $X\sim{\cal N}\left(0,\frac{1}{2}\right)$.
@@
 Since the range of $X$ is unbounded, we truncate its range within [b,b][-b,b] for a given parameter $b>0$.
-Accordingly, the range of the CDF in the positive region should be modified as $\left[\frac{1}{2},F_{X}(b)\right]$.
-Thus, the tentative value $\hat{a}_{k}$, which corresponds to the point $a_{k}$ before energy normalization, is determined by the following equality:
+Accordingly, the range of the CDF in the positive amplitude region should be modified as $\left[\frac{1}{2},F_{X}(b)\right]$.
+Then, the tentative value $\hat{a}_{k}$, which corresponds to the point $a_{k}$ before energy normalization, is determined by the following equality:
 $F_{X}(\hat{a}_{k})=\frac{1}{2}+l_{k}\frac{F_{X}(b)-\frac{1}{2}}{M},$ (2)
@@
 $l_{k}=2k-1,\quad\,\text{ for }\,k\in\mathcal{M}_{+}$ (3)
-is the integer value unique to the $k$th constellation point.
-Let us define $\rho$ as
-$\rho\triangleq\operatorname{erf}(b),\quad\rho\in(0,1].$ (4)
-We may then express $\hat{a}_{k}$ in (2) using (1) as
+is the integer value unique to the $k$th constellation point. Let us define $\rho$ as
+$\rho\triangleq\operatorname{erf}(b),\quad\rho\in(0,1]$ (4)
+such that the CDF in the positive range of amplitude reduces to $\left[\frac{1}{2},\frac{1+\rho}{2}\right]$ according to (1). We may then express $\hat{a}_{k}$ in (2) using (1) as
 $\hat{a}_{k}=\phi_{M}(l_{k};\rho)\triangleq\operatorname{erf}^{-1}\left(\rho\frac{l_{k}}{M}\right),$ (5)
@@
 Finally, scaling the magnitude of $\hat{a}_{k}$ in accordance with the energy constraint leads to the desired constellation point $a_{k}$:
-$a_{k}=\sqrt{\frac{E_{s}}{2}}\frac{\hat{a}_{k}}{\sqrt{\frac{2}{M}\sum_{k=1}^{M/2}\hat{a}^{2}_{k}}},$ (6)
-where $E_{s}$ is the energy in terms of QAM symbol.
-This design algorithm uniquely determines the resulting constellation using a single parameter $\rho\in(0,1]$ for a given modulation order $M$.
-When $\rho=1$ (i.e., without truncation), the resulting constellation coincides with the Gaussian-like PAM proposed in [21].
+$a_{k}=\sqrt{\frac{E_{s}}{2}}\frac{\hat{a}_{k}}{\sqrt{\frac{2}{M}\sum_{\ell=1}^{M/2}\hat{a}^{2}_{\ell}}},$ (6)
+where $E_{s}$ denotes the average transmit energy per complex-valued symbol. In our subsequent numerical calculation of the constellation points, we set $E_{s}=1$ for simplicity.
+The above design algorithm uniquely determines the non-uniform PAM constellation using a single parameter $\rho\in(0,1]$ for a given modulation order $M$.
+When $\rho=1$ (i.e., without truncation), the resulting constellation coincides with the Gaussian-like PAM proposed in [boutros2018geometric].
 In contrast, we may express the Maclaurin series of $F_{X}(x)$ up to the first order as
@@
 which indicates that $F_{X}(x)$ linearly increases (i.e., $X$ becomes a uniform distribution) for $X\in[-b,b]$ in the limit of $b\to 0$ ($\rho\to 0$).
-For this reason, we denote $\rho=0$ as the standard PAM in what follows.
-Figure 1: Constellation point selection based on equiprobable partitioning of the Gaussian CDF ($\rho=1$) with $M=32$. Due to the symmetry, only the positive region is shown.
-The relationship between the CDF values and the selected constellation points $\{\hat{a}_{k}\}$ in the case of $M=32$ and $\rho=1$ (i.e., without range truncation) is illustrated in Fig.
-1.
-As shown in the figure, the values $\phi_{M}(2(k-1);\rho)$ and $\phi_{M}(2k;\rho)$ serve as the boundary of the $k$th equiprobable sub-region, which are depicted as the corresponding gray dashed lines.
-Each sub-region is assigned a single bit label, and the representative constellation point $\hat{a}_{k}$ is then selected to match the midpoint CDF value (i.e., $\phi_{M}(l_{k};\rho)$) of its corresponding region.
+For this reason, we denote $\rho=0$ as the standard uniform PAM in what follows.
+(a) $\rho=1.0$
+(b) $\rho=0.5$
+Figure 1: Constellation point selection based on equiprobable partitioning of the Gaussian CDF with $M=32$.
+Due to the symmetry, only the positive region is shown.
+(a) is the case with $\rho=1.0$ (i.e., without truncation), while (b) corresponds to the case with $\rho=0.5$
+The relationship between the CDF values and the selected constellation points $\{\hat{a}_{k}\}$ for the case of $M=32$ and $\rho=1$ (i.e., without truncation) is illustrated in Fig.
+1(a).
+The values $\phi_{M}(2(k-1);\rho)$ and $\phi_{M}(2k;\rho)$ serve as the boundaries of the $k$th equiprobable subregion, which are depicted as the corresponding gray dashed lines.
+Each subregion is assigned a single binary label ${\bf b}\in\mathbb{F}_{2}^{m}$, and the representative constellation point $\hat{a}_{k}$ is then selected to match the midpoint CDF value (i.e., $\phi_{M}(l_{k};\rho)$) of the corresponding region.
+On the other hand, the case of $\rho=0.5$ is shown in Fig.
+1(b), where the effective range of the original Gaussian CDF is now truncated to $\left[\frac{1}{2},\frac{1+\rho}{2}\right]=[0.5,0.75]$, as indicated by the red dashed line.
+Within this truncated range, we observe that the CDF increases almost linearly, and thus the resulting constellation points become more uniformly spaced, approaching those of the conventional uniform PAM constellation.
 II-B Constellation Point Merging Process
-Since the optimal constellation cardinality that maximizes the BMI varies depending on the channel SNR, we consider dynamically adjusting the number of constellation points by incorporating a constellation point merging strategy into the truncated Gaussian-based design described above.
-We define $N$ as the number of distinct points in $M$-ary PAM symbol with $N\leq M$, which means that a PAM constellation with $M$ distinct labels has only $N$ distinct constellation points.
+In coded modulation systems that do not utilize bit-wise interleaving, the achievable information rate, referred to as the constellation-constrained capacity, consistently increases with higher-order modulations.
+In BICM systems, the sum of the bit-wise mutual information (BMI) serves as the achievable information rate, and the constellation size that maximizes the total BMI depends on the channel SNR, indicating that high-order constellations may become suboptimal in low-SNR regimes [caire1998bit, zollner2013optimization].
+Therefore, we next consider dynamically adjusting the number of constellation points by incorporating a strategy of merging constellation points into the truncated Gaussian-based geometrical design described above.
+We define $N$ as the number of distinct points in an $M$-ary PAM symbol with $N\leq M$, which indicates that a PAM constellation with $M$ distinct labels has only $N$ distinct constellation points.
 We denote this constellation as $(M,N)$-PAM in this work.
@@
 Therefore, for an initial modulation order of $M$, the procedure can be executed a maximum of M/21M/2-1 times in principle, ultimately leading to the standard BPSK constellation (when $N=2$).
-Each merging operation combines the corresponding sub-regions into a new, unified sub-region, and the representative constellation point is placed to match the midpoint CDF value of this modified sub-region.
+Each merging operation combines the corresponding subregions into a new, unified subregion, and the representative constellation point is placed to match the midpoint CDF value of this modified subregion.
 Now, suppose that for a given parameter $\rho$, the constellation points $\hat{a}_{i}=\phi_{M}(l_{i};\rho)$ and $\hat{a}_{j}=\phi_{M}(l_{j};\rho)$ (with $i\neq j$) have the minimum Euclidean distance at a certain merging step.
-Then, these two points are merged into a single point by updating their indices and amplitude values as222 Since the initial indices $\{l_{k}\}$ are set to be positive odd numbers according to (3), the first merging step always applies to the two positive odd numbers.
+Then, these two points are merged into a single point by updating their indices and amplitude values as [^3] Since the initial indices $\{l_{k}\}$ are set to be positive odd numbers according to (3), the first merging step always applies to the two positive odd numbers.
 As a result, the merged index $l^{\prime}$ is guaranteed to be a natural number.
 $\hat{a}_{i}=\hat{a}_{j}:=\phi_{M}\left(l^{\prime};\rho\right),\quad l^{\prime}:=\frac{l_{i}+l_{j}}{2}.$ (7)
-After this process, the multiple bit labels originally associated with $\hat{a}_{i}$ and $\hat{a}_{j}$ are mapped to the identical signal amplitude.
-The proposed constellation shaping algorithm for a given pair of initial modulation order $M$ and the number of distinct points $N$ after the merging process is summarized in Algorithm 1, where $\mathcal{L}$ corresponds to the set of the constellation indices in the positive region that have unique constellation point locations.
+After this process, the multiple binary labels originally associated with $\hat{a}_{i}$ and $\hat{a}_{j}$ are mapped to the identical signal amplitude.
+The proposed shaping algorithm, which is summarized in Algorithm 1, applies to a specific pair of initial modulation order $M$ and the number of distinct points $N$ after the merging process, with $\mathcal{L}$ representing the set of the constellation indices in the positive amplitude region that have unique constellation point locations.
 The algorithm thus terminates when $|\mathcal{L}|=N/2$.
-Figure 2: Constellation point selection based on non-equiprobable partitioning of the Gaussian CDF ($\rho=1$) with $M=32$ and $N=10$.
+In what follows, a set of the resulting distinct PAM constellation points located in both positive and negative regions, after normalization through (6), will be denoted by ${\cal A}\left(M,N,\rho\right)$, which is uniquely determined by the two parameters, $N$ and $\rho$, for a given modulation order $M$.
+Furthermore, since each binary label is converted to a corresponding PAM symbol, we define the symbol mapping function as
+$\varphi:{\mathbb{F}}_{2}^{m}\to{\cal A}\left(M,N,\rho\right)$ (8)
+such that $\varphi(\bf b)$ represents the PAM symbol that has the binary label of length $m$ specified by $\bf b$}.
+We note that if $N<M$ the mapping function $\varphi$ will be characterized by many-to-one mapping.
+Figure 2: Constellation point selection based on non-equiprobable partitioning of the Gaussian CDF ($\rho=1.0$) with $M=32$ and $N=10$.
 Algorithm 1 Construction of $(M,N)$-PAM
@@
 TABLE I: The $l_{k}$ indices of constellation points $\hat{\mathcal{A}}_{+}$ for the $(32,N)$-PAM with $\rho=1$ (no truncation).
-The second row indicates the bit labels.
+The second row indicates the binary labels.
 The initial amplitude values are shown in the third row, while the values after the merging process are shown in brackets.
-The number in the leftmost column represents that of distinct points $N$.
+The number in the leftmost column represents NN (distinct points).
 $k$ 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16
-bit label 00000 00001 00011 00010 00110 00111 00101 00100 01100 01101 01111 01110 01010 01011 01001 01000
+label 00000 00001 00011 00010 00110 00111 00101 00100 01100 01101 01111 01110 01010 01011 01001 01000
 $\hat{a}_{k}$ .028 .083 .139 .196 .255 .315 .377 .443 .512 .587 .669 .762 .870 1.003 1.185 1.523
 32 1 3 5 7 9 11 13 15 17 19 21 23 25 27 29 31
-30 2 [.055] 5 7 9 11 13 15 17 19 21 23 25 27 29 31
-28 2 6 [.168] 9 11 13 15 17 19 21 23 25 27 29 31
-26 2 6 10 [.284] 13 15 17 19 21 23 25 27 29 31
-24 2 6 10 14 [.410] 17 19 21 23 25 27 29 31
-22 2 6 10 14 18 [.549] 21 23 25 27 29 31
-20 2 6 10 14 18 22 [.714] 25 27 29 31
-18 4 [.111] 10 14 18 22 25 27 29 31
-16 4 12 [.346] 18 22 25 27 29 31
-14 4 12 18 22 26 [.932] 29 31
-12 4 12 20 [.627] 26 29 31
-10 8 [.225] 20 26 29 31
-8 8 20 26 30 [1.317]
-6 8 20 28 [1.085]
-4 8 24 [.813]
-2 16 [.477]
-The integer representations $\{l_{k}\}$ of the constellation points for the positive region $\hat{\mathcal{A}}_{+}$ obtained by Algorithm 1 are listed in TABLE I in the case of $(32,N)$-PAM starting from $N=32$ to $2$, where the initial constellation points are derived from the original Gaussian distribution (i.e., $\rho=1$).
-The bit labels are based on BRGC.
-Therefore, the bit label of the negative region $a_{k}$ is identical to that of $a_{k}$ with MSB replaced by $1$.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+The integer representations $\{l_{k}\}$ of the constellation points for the positive region $\hat{\mathcal{A}}_{+}$ generated by Algorithm 1 are listed in Table I in the case of $(32,N)$-PAM starting from $N=32$ to $2$, where the initial constellation points are derived from the original Gaussian distribution (i.e., $\rho=1$).
+Since the binary labels are based on BRGC, the bit labels of the negative region $a_{k}$ are identical to that of $a_{k}$ with the MSB replaced by $1$.
 Focusing on the first merging step from $N=32$ to $30$, the initial constellation points near the origin (e.g., $\hat{a}_{1},\hat{a}_{2},\hat{a}_{3},\hat{a}_{4}$) exhibit very small Euclidean distances.
-When the adjacent point pair $\hat{a}_{1}$ and $\hat{a}_{2}$ is merged, they are modified into a single signal point but with two distinct bit labels.
-Then, these merged symbols become indistinguishable from each other, but they become more separable from the remaining symbols (e.g., $\hat{a}_{3},\hat{a}_{4}$).
-The conventional Gray labeling ensures that the interference from adjacent symbols results in only a single-bit error.
-However, in high-order constellations under low SNR conditions, the interference from non-adjacent symbols become likely as well, resulting in multiple-bit errors.
-The proposed constellation merging strategy effectively mitigates this issue by adaptively controlling the balance of density in terms of geometry and probability.
+When the adjacent point pair $\hat{a}_{1}$ and $\hat{a}_{2}$ is merged, they are modified into a single constellation point that is assigned two distinct binary labels.
+Then, these merged symbols become indistinguishable from each other, while they become more separable from the remaining symbols (e.g., $\hat{a}_{3},\hat{a}_{4}$).
+As a result, the constellation behaves similarly to a lower-order constellation with fewer but more reliable symbols.
 Fig.
 2 illustrates the constellation point selection process based on the proposed merging scheme applied to the resulting constellation of Fig.
-1, with the number of distinct points set to $N=10$.
-As shown in the figure, the original equiprobable sub-regions (gray dashed lines) are merged into larger sub-regions, depicted as blue dashed lines.
+1(a), with the number of distinct points set to $N=10$.
+As shown in the figure, the original equiprobable subregions (gray dashed lines) are merged into larger subregions, depicted as blue dashed lines.
 Each representative constellation point is then selected in the same manner as that in Fig.
-1, but based on the midpoint CDF value of the modified sub-region.
-This merging process corresponds to a non-equiprobable partitioning of the reference Gaussian distribution, according to the number of bit labels assigned to each constellation point.
-We note that constellation point merging or many-to-one mapping in [26, 28, 29] refers to a low-complexity technique for implementing probabilistic shaping, where redundant bit labels are assigned to control the probability mass function (PMF) of the constellation to achieve shaping gain.
-In contrast, the proposed approach utilizes constellation point merging as an internal tool of geometric constellation design for the purpose of avoiding excessive density in the constellation.
-Therefore, this modification does not affect the complexity of the transmitter.
+1, but based on the midpoint CDF value of the modified subregion.
+This merging process corresponds to a non-equiprobable partitioning of the reference Gaussian distribution, according to the number of binary labels assigned to each constellation point.
+Considering the fact that the resulting constellation exhibits a non-uniform probability mass function (PMF), the proposed design can be viewed as a hybrid of geometric and probabilistic shaping.
+However, the non-uniformity in symbol probabilities arises from the rearrangement (i.e., geometrical design) of the Gray-labeled $M$ symbols, where multiple symbols are merged to have an identical amplitude.
+Thus, unlike conventional probabilistic shaping schemes that require a distribution matcher, the proposed shaping scheme imposes only the update of the bit-to-symbol mapping table in a PAM-based BICM system with the same modulation order $M$.
 II-D Comparison With NUC of ATSC 3.0
 Figure 3: Comparison between the proposed QAM consisting of two (32,20)-PAM constellations (shown as green circles) and the equivalent one-dimensional NUC adopted in ATSC 3.0 (shown as blue diamonds).
-Both constellations are optimized for BICM with an information rate of 8.333 bits per symbol (code rate of $3/5$).
-The proposed QAM constellation points consisting of two $(32,20)$-PAM constellations where the parameter $\rho$ is optimized to maximize the BMI at the information rate corresponding to the code rate of $3/5$ (discussed in the next section) are depicted as green circles in Fig.
-3.
-For comparison, those of the equivalent $1024$-point NUC adopted in ATSC 3.0 are plotted as blue diamonds.
-While the ATSC NUC individually optimizes all $32$ points per each dimension, many points around the origin are almost overlapping, resembling a smaller number of distinct points.
-Interestingly, the proposed constellation, systematically derived by Algorithm 1 and governed by only the two design parameters $N$ and $\rho$, exhibits a remarkably similar configuration to the NUC of ATSC 3.0.
-The performance analysis and optimization of these parameters leading to the above result are presented in the following section.
+Both constellations are optimized for BICM with an information rate of 6.0 bits per symbol (code rate of $3/5$).
+The size of the circles reflects their probability of occurrence.
+The proposed NENU constellation consisting of two $(32,20)$-PAM constellations is depicted using green circles in Fig.
+3, where symbols with higher probabilities are plotted with larger marks.
+The parameter $\rho$ is optimized based on the BMI for a target spectral efficiency of $6.0$ bits per symbol, corresponding to a code rate of $3/5$.
+For comparison, the equivalent $1024$-ary one-dimensional NUC adopted in ATSC 3.0 is depicted with blue diamonds.
+While the ATSC NUC individually optimizes all PAM amplitudes, many points near the origin are almost overlapping, resembling a smaller number of distinct points.
+The proposed constellation, which is systematically derived by Algorithm 1 and governed by only the two design parameters $N$ and $\rho$, exhibits a remarkably similar configuration to the NUC of ATSC 3.0.
+The performance analysis and optimization of these parameters leading to the above result are described in the following section.
 III Performance Analysis
@@
 We compare the results with other constellations found through numerical optimizations, which require significantly higher complexity than the proposed scheme.
-From Section II-B, for a given modulation order $M$, the proposed constellation is specified by the two parameters, $N$ and $\rho$.
-We denote the resulting constellation points located in both positive and negative regions by ${\cal A}\left(M,N,\rho\right)$ in what follows.
-III-A Total BMI
-To identify the constellations that are expected to perform well, it is necessary to evaluate their achievable information rates.
-In BICM systems, the achievable information rate is defined by the BMI [1].
-For a given $M$-ary PAM constellation $\cal A$}, the total BMI at a specified channel SNR $\gamma$ is given by
-$C\left(\mathcal{A},\gamma\right)=n-\sum_{i=0}^{n-1}{\mathbb{E}_{Y,b}\left[\log_{2}\frac{\sum_{x\in\mathcal{A}}p_{Y\mid X}(Y\,|\,x)}{\sum_{x\in\mathcal{A}_{b}^{(i)}}p_{Y\,|\,X}(Y\,|\,x)}\right]},$ (8)
-measured in bits per dimension, where $n=\log_{2}M$ is the total number of binary channels and $\mathcal{A}_{b}^{(i)}\subset{\cal A}$ denotes the set of constellation points in the $i$th binary channel with the bit specified by $b\in\mathbb{F}_{2}$.
-The conditional probability density function (PDF) $p_{Y|X}(y\,|\,x)$ is defined as
+III-A Bit-wise Mutual Information
+Let ${\mathcal{B}}_{b}^{(i)}\subset\mathbb{F}_{2}^{m}$ denote a set of binary labels with its $i$th bit specified by $b\in\mathbb{F}_{2}$.
+Let $\cal A$} denote a set of $M$-ary PAM constellation points that are mapped from the binary label through (8).
+The BMI of the $i$th binary channel at a specified channel SNR $\gamma$ over an AWGN channel is then expressed as [szczecinski2015bit, caire1998bit]
+$C_{i}\left(\mathcal{A},\gamma\right)=1-\sum_{b\in\mathbb{F}_{2}}\sum_{{\bf b}\in\mathcal{B}_{b}^{(i)}}P({\bf b})\,$
+$\times\int_{-\infty}^{\infty}p_{Y\mid X}(y\mid\varphi({\bf b}))\log_{2}\frac{\sum_{{\bf b}^{\prime}\in\mathbb{F}_{2}^{m}}p_{Y\mid X}(y\,|\,\varphi({\bf b}^{\prime}))}{\sum_{{\bf b}^{\prime\prime}\in\mathcal{B}_{b}^{(i)}}p_{Y\mid X}(y\,|\,\varphi({\bf b}^{\prime\prime}))}dy,$ (9)
+where $P({\bf b})=\frac{1}{M}$ (i.e., the input binary labels are equiprobable) and the conditional probability density function (PDF) $p_{Y|X}(y\,|\,x)$ is defined as
 $p_{Y|X}(y\,|\,x)=\frac{1}{\sqrt{\pi N_{0}}}e^{-\frac{(y-x)^{2}}{N_{0}}},$ (10)
-where $N_{0}$ represents the power spectral density of AWGN, which is related to $\gamma$ by $\gamma=E_{s}/N_{0}$.
-III-B Parameter Optimization
-For a given pair of modulation order $M$ and operating SNR $\gamma$, the optimization problem in our NENU constellation is formulated as
+with $N_{0}$ representing the power spectral density of the AWGN, which is related to $\gamma$ by $\gamma=E_{s}/N_{0}$.
+Note that The above integral can be efficiently evaluated using Gaussian–Hermite quadrature (GHQ) [alvarado2018achievable].
+Accordingly, the achievable information rate of a BICM system is defined as the sum of the BMIs over all binary channels:
+$C\left(\mathcal{A},\gamma\right)=\sum_{i=0}^{m-1}C_{i}(\mathcal{A},\gamma),$ (11)
+measured in bits per dimension.
+TABLE II: BMI of individual binary channels for $(32,N)$-PAM with $\rho=1$. The bit $b^{(0)}$ corresponds to the MSB.
+NN (SNR) $b^{(0)}$ $b^{(1)}$ $b^{(2)}$ $b^{(3)}$ $b^{(4)}$ Total
+32 (10 dB) 0.6940 0.5163 0.2291 0.0741 0.0268 1.5403
+10 (10 dB) 0.7518 0.5915 0.1739 0.0695 0.0262 1.6128
+32 (20 dB) 0.9013 0.8453 0.7193 0.4800 0.2239 3.1699
+10 (20 dB) 0.9990 0.9953 0.4812 0.2241 0.1158 2.8154
+[…]
+III-B Effect of Constellation Point Merging on BMI
+To elucidate the validity of the proposed merging operation, we first demonstrate its impact on the BMI of individual binary channels.
+As an example of $M=32$ (corresponding to five binary channels), Table II lists the BMI values of individual binary channels for $(32,32)$-PAM and $(32,10)$-PAM ($N=32$ and $N=10$), each evaluated at SNRs of $10$ dB and $20$ dB, where $b^{(0)}$ and $b^{(4)}$ correspond to the MSB and the LSB, respectively.
+It is observed in both SNR cases that the BMIs of the lower-order bits, such as $b^{(3)}$ and $b^{(4)}$, are reduced by the merging process, while those of the higher-order bits, $b^{(0)}$ and $b^{(1)}$, are increased.
+Therefore, the proposed merging operation enhances the reliability of the higher-order bits at the expense of that of the lower-order bits.
+In the case of lower SNR ($10$ dB), in particular, the lower-order bits are inherently unreliable and contribute only marginally to the total BMI.
+As a result, the loss in their BMIs due to merging is relatively small, whereas the enhanced separability among constellation points successfully increases the BMIs of the higher-order bits, leading to an overall improvement in the total BMI.
+III-C Loss From the Shannon Limit
+Figure 4: Loss in achievable information rate from the Shannon limit for the BICM systems using the proposed NENU constellations with various modulation orders.
+We evaluate the gain achieved by the proposed shaping scheme when its parameters are properly optimized.
+For a given pair of modulation order $M$ and operating SNR $\gamma$, the optimization problem for our NENU constellation is formulated as
 $\max_{N,\rho}$ $C\Bigl({\cal A}\left(M,N,\rho\right),\gamma\Bigr),$ (12)
@@
 $\rho\in[0,1].$
-By evaluating the BMI over all parameter pairs of $(N,\rho)$, the optimal constellation can be identified.
-(In practice, we may apply the quantized value of $\rho$ with the step size of $0.01$.) Note that when the channel is modeled as AWGN, the expectation over the Gaussian noise in the BMI calculation (8) can be efficiently computed using Gaussian–Hermite quadrature (GHQ) [35].
-III-C Loss From the Shannon Limit
-Figure 4: Loss in achievable information rate from the Shannon limit for the BICM systems using the proposed constellations with various modulation orders.
-The gain achieved by constellation shaping is typically assessed by the difference between the achievable information rate and the Shannon capacity, known as the loss from the Shannon limit [36].
+In this work, constellations with varying numbers of distinct points $N\in\{2,4,\ldots,M\}$ are generated for each quantized value of $\rho$, using a step size of $0.01$, in accordance with the proposed merging principle.
+The optimal constellation is determined by evaluating the BMI in (11) across all candidate parameter pairs $(N,\rho)$.
+It is important to note that if $\rho$ is fixed for a specific target SNR and $N$ is subsequently adjusted to enhance the BMI, the optimization process would be simplified.
+However, this approach may lead to a suboptimal constellation.
+The gain achieved by constellation shaping is typically assessed by the difference between the achievable information rate and the Shannon limit, known as the loss from the Shannon limit [montorsi2016design].
 When the two identical $M$-ary PAM constellations form $M^{2}$-ary QAM, the loss from the Shannon limit is defined as a function of SNR $\gamma$ by
@@
 Each plotted point is obtained by selecting the optimal parameter set $\left(N,\rho\right)$ at a given SNR.
-As a comparison, the minimum loss achieved by the uniform $M^{2}$-QAM with the optimized modulation order is indicated as a gray dashed line.
-At low SNRs, the performance of the proposed shaping scheme coincides with that of conventional QAM, due to the optimality of low-order constellations such as QPSK and uniform 16-QAM in this SNR regime.
-However, as the SNR increases, the gap between the BMI of standard QAM and the Shannon limit becomes noticeable.
-In contrast, the proposed NENU constellation consistently benefits from higher modulation orders across all SNRs.
-This indicates that, due to its flexible control over the number of signal points, properly designed NENU constellations successfully avoid excessive geometric density that would otherwise degrade the BMI.
-In fact, the proposed NENU constellation enhances the BMI because of an effect similar to probabilistic shaping.
+As a comparison, the minimum loss achieved by the uniform $M^{2}$-QAM with the optimal modulation order is indicated as a gray dashed line.
+At low SNRs, the performance of the proposed shaping scheme coincides with that of conventional QAM, due to the optimality of low-order constellations such as QPSK and standard uniform 16-QAM in this SNR regime.
+However, as the SNR increases, the gap between the BMI of standard QAM and the Shannon limit becomes increasingly pronounced.
+On the other hand, the proposed NENU constellation achieves a minimal loss from the Shannon limit, particularly in the high-SNR regimes, by exploiting higher modulation orders.
+Specifically, losses of less than $0.1$ bits per symbol are observed for modulation orders of 16 38416\,384 and above.
+These remarkable gains are attributed to the high-resolution Gaussian approximation enabled by the huge number of constellation points.
+Moreover, higher-order NENU constellations consistently outperform lower-order ones across the entire SNR range.
+This confirms that the proposed merging scheme adaptively reduces the effective constellation size of high-order NENU constellations in the low-SNR regime.
 III-D Comparison With Other Geometric Shaping Schemes
 Figure 5: Loss from the Shannon limit for the BICM systems with several different modulation schemes.
-We next compare the proposed NENU constellations with other constellations.
-For this purpose, we fix the modulation order as $M^{2}=1024$ in what follows.
-We consider the following numerically optimized constellations for comparison:
+We next compare the proposed NENU constellations with other representative modulation schemes. The following constellations are considered for comparison:
+  • The uniform $M^{2}$-ary QAM with $M^{2}\in\{64,256,1024\}$
+  • The proposed one-dimensional shaping without constellation point merging, i.e., ${\cal A}\left(M,M,\rho\right)$ with $M^{2}\in\{64,256,1024\}$ (indicated as shaped QAM).
+  • The proposed NENU constellation ${\cal A}\left(32,N,\rho\right)$.
   • The one-dimensional (1D) $1024$-NUC of ATSC 3.0.
-  • The two-dimensional (2D) $1024$-NUC developed in [7].
-  • The uniform $M^{2}$-ary QAM with $M^{2}\in\{64,256,1024\}$
-  • The one-dimensional shaping without constellation merging, i.e., ${\cal A}\left(M,M,\rho\right)$ with $M^{2}\in\{64,256,1024\}$ (indicated as shaped QAM).
-  • The proposed one-dimensional NENU constellation ${\cal A}\left(32,N,\rho\right)$.
-The results are compared in Fig.
-5.
-We first observe that the optimal modulation order in BICM is contingent upon the channel SNR.
-The shaped QAM without constellation merging has been confirmed to demonstrate improved performance compared to uniform QAM [23].
-However, due to the constraint of constellation sizes to powers of two, these schemes exhibit significant gaps from the Shannon limit at specific SNR regions, particularly around 15 dB and 20 dB.
-In the case of the proposed shaping scheme, these gaps can be maintained at a minimal level across a wide range of SNR.
-This suggests that employing shaping with a flexible constellation size is advantageous for rate-adaptive BICM design.
-The gain achieved through the proposed shaping scheme is comparable to that of the one-dimensional NUC in ATSC 3.0 across all the SNRs, even though the latter is obtained by a more complex optimization process.
-When comparing the results with the two-dimensional constellation of [7], performance gaps of approximately $0.08$ bits per symbol are observed, highlighting the limitations inherent in one-dimensional constellation design.
-However, one-dimensional modulation schemes offer practical advantages in very high data-rate regimes, since the demapping process for two-dimensional constellations requires a considerable increase in computational cost with higher modulation orders, as is discussed in the next subsection.
+  • The two-dimensional (2D) $1024$-NUC developed in [sillekens2022high].
+The corresponding loss from the Shannon limit is compared in Fig.
+5, where all shaping schemes other than standard uniform QAM are optimized at each plotted SNR according to their respective design or optimization algorithms.
+We begin with a brief discussion of the design complexity in terms of the number of real-valued variables to be optimized (i.e., the degrees of freedom).
+Let $m^{\prime}=2m=\log_{2}M^{2}$ denote the number of binary channels (i.e., the maximum spectral efficiency in bits per symbol).
+For the one-dimensional NUC in ATSC, $2^{m^{\prime}/2-1}=2^{m-1}$ real-valued PAM amplitudes need to be optimized under an origin-symmetry constraint.
+On the other hand, the two-dimensional NUC in [sillekens2022high] is designed directly in the complex plane under a quadrant symmetry.
+In this case, $2^{m^{\prime}-2}=2^{2(m-1)}$ complex-valued constellation points in a single quadrant must be optimized, which corresponds to $2^{2m-1}$ real-valued variables for the I/Q components.
+In contrast, for the proposed NENU constellation, all PAM amplitudes are optimized using only two real-valued parameters irrespective of $m$, highlighting its scalability for high-spectral-efficiency regimes.
+As shown in Fig.
+5, the analysis of the achievable rate reveals that the optimal modulation order in BICM is influenced by the channel SNR.
+The implementation of shaped QAM without merging constellation points has been validated to provide better performance compared to standard uniform QAM [kurihara2024design].
+However, the limitation of constellation sizes being restricted to powers of two results in notable gaps from the Shannon limit at certain SNR ranges, specifically around 15 dB and 20 dB.
+In contrast, the proposed shaping scheme maintains these gaps at a minimal level over a broad range of SNRs.
+This indicates that using shaping with the ability to flexibly adjust both the constellation size and its distribution is beneficial for a rate-adaptive BICM design.
+The performance gain from the proposed shaping scheme is comparable to that of the one-dimensional NUC in ATSC 3.0 across all SNR ranges, despite the latter being achieved through a more complex optimization process for a specific target SNR or information rate.
+When comparing the results with the two-dimensional constellation of [sillekens2022high], performance gaps of approximately $0.08$ bits per symbol are observed, which highlights the inherent limitations of one-dimensional constellation design.
+However, one-dimensional modulation schemes offer practical advantages in very high data-rate regimes, since the demapping process for two-dimensional constellations requires a substantial increase in computational cost with higher modulation orders, as is discussed in the next subsection.
 III-E Demapping Complexity
+In BICM systems, soft demapping is carried out by computing the bit log-likelihood ratios (LLRs) for each binary channel, which dominates the computational complexity at the receiver.
+The required computational effort strongly depends on the geometric structure of the underlying constellation and becomes increasingly critical as the modulation order increases.
+Accordingly, we analyze the LLR computation complexity for general one-dimensional and two-dimensional constellations and then discuss the corresponding complexity of the proposed NENU constellation.
+Let us assume that the transmitted symbols are selected from a set of specific complex-valued (i.e., two-dimensional) constellation $\mathcal{X}$, with the number of points represented by $|\mathcal{X}|$.
+The LLR value for the $i$th binary channel is computed as
+$L_{i}$ $=\log\left[\frac{\sum_{x\in\mathcal{X}_{0}^{(i)}}p(x)\exp\left(-\frac{|y-x|^{2}}{N_{0}}\right)}{\sum_{x\in\mathcal{X}_{1}^{(i)}}p(x)\exp\left(-\frac{|y-x|^{2}}{N_{0}}\right)}\right]$
+$\approx\frac{1}{N_{0}}\left(\min_{x\in\mathcal{X}_{1}^{\left(i\right)}}\left\{|y-x|^{2}\right\}-\min_{x\in\mathcal{X}_{0}^{\left(i\right)}}\left\{|y-x|^{2}\right\}\right),$ (14)
+where $y\in{\mathbb{C}}$ denotes the received complex-valued symbol and $\mathcal{X}_{b}^{(i)}\subset\mathcal{X}$ represents the set of constellation points in the $i$th binary channel with the bit specified by $b\in\mathbb{F}_{2}$.
+On the other hand, for a one-dimensional constellation, where $\mathcal{X}$ is composed of two identical and orthogonal PAM constellations $\mathcal{A}$, the LLR value can be expressed as
+$L_{i}$ $\approx\frac{1}{N_{0}}\left(\min_{x\in\mathcal{A}_{1}^{\left(i\right)}}\left\{(y-x)^{2}\right\}-\min_{x\in\mathcal{A}_{0}^{\left(i\right)}}\left\{(y-x)^{2}\right\}\right),$ (15)
+where $y\in{\mathbb{R}}$ is the received real-valued symbol, $\mathcal{A}_{b}^{(i)}\subset\mathcal{A}$ denotes the set of constellation points in the $i$th binary channel with the bit specified by $b\in\mathbb{F}_{2}$, and $x$ is thus chosen from a one-dimensional (i.e., real-valued) constellation.
+We focus only on the number of required multiplication operations, since the complexity associated with comparison and addition operations is negligible compared to that of multiplication operations.
+In the case of a two-dimensional constellation, from (14) it is necessary to compute the squared Euclidean distances between the received symbol and each of the $|\mathcal{X}|=M^{2}$ candidate constellation points.
+This computation entails performing $2|\mathcal{X}|$ real multiplications.
+Meanwhile, in the case of a one-dimensional constellation, calculating (15) requires $|\mathcal{A}|$ real multiplications.
+The square root of the cardinality of a two-dimensional constellation, $|\mathcal{A}|=M$, represents the corresponding cardinality in this case.
+Therefore, as $M$ increases to support higher spectral efficiency, the use of a one-dimensional constellation should be critical from an implementation complexity perspective.
+This reasoning may elucidate the rationale behind the adoption of one-dimensional constellations for high-order scenarios in practical wireless communication systems such as ATSC 3.0.
 Figure 6: The optimal values of constellation size $N$ and design parameter $\rho$ according to the channel SNR ($M=32$).
-For efficient decoding, it is necessary to compute bit log-likelihood ratio (LLR) and this dominates the computational complexity at the receiver especially when irregular constellations are adopted.
-In what follows, we discuss the demapping complexity for general QAM constellations  [^3] There have been several attempts to simplify the soft-output demappers for both one-dimensional and two-dimensional non-uniform constellations [37, 38, 39, 40], but they rely on specific geometric structures of constellations.
-.
-Let us assume that the transmitted symbols are selected from a set of specific complex-valued (i.e., two-dimensional) constellation set $\mathcal{A}$, with the number of points represented by $|\mathcal{A}|$.
-The LLR value for the $i$th bit level is computed as
-$L_{i}$ $=\log\left[\frac{\sum_{x\in\mathcal{A}_{0}^{(i)}}\exp\left(-\frac{|y-x|^{2}}{N_{0}}\right)}{\sum_{x\in\mathcal{A}_{1}^{(i)}}\exp\left(-\frac{|y-x|^{2}}{N_{0}}\right)}\right]$
-$\approx\frac{1}{N_{0}}\left(\min_{x\in\mathcal{A}_{1}^{\left(i\right)}}\left\{|y-x|^{2}\right\}-\min_{x\in\mathcal{A}_{0}^{\left(i\right)}}\left\{|y-x|^{2}\right\}\right),$ (12)
-where $y\in{\mathbb{C}}$ denotes the received complex-valued symbol.
-We focus only on the number of required multiplication operations as the complexity associated with comparison and addition operations is considered negligible in comparison to that of multiplication operations.
-On the other hand, if $\mathcal{A}$ is a set of one-dimensional constellation points, the LLR value can be expressed as
-$L_{i}$ $\approx\frac{1}{N_{0}}\left(\min_{x\in\mathcal{A}_{1}^{\left(i\right)}}\left\{(y-x)^{2}\right\}-\min_{x\in\mathcal{A}_{0}^{\left(i\right)}}\left\{(y-x)^{2}\right\}\right),$ (13)
-where $y\in{\mathbb{R}}$ is the received real-valued symbol and $x$ is chosen from one-dimensional (i.e., real-valued) constellation.
-In the case of a two-dimensional constellation, from (12) it is necessary to compute the squared Euclidean distances between the received symbol and each of the $|\mathcal{A}|$ constellation points.
-This computation entails performing $2|\mathcal{A}|$ real multiplications.
-Furthermore, the cardinality of the two-dimensional constellations is $|\mathcal{A}|=M^{2}$.
-On the other hand, in the case of a one-dimensional constellation, calculating (13) requires $|\mathcal{A}|$ real multiplications, and the corresponding cardinality is $|\mathcal{A}|=M$.
-Therefore, as $M$ increases for higher spectral efficiency, the use of a one-dimensional constellation should be critical in view of practical implementation.
-This reasoning may elucidate the rationale behind the adoption of one-dimensional constellations for high-order scenarios in practical wireless communication systems such as ATSC 3.0, despite its lower achievable gain.
-The proposed NENU constellation also offers a more efficient demapping implementation due to its reduced constellation size from $M$, since the demapping complexity is proportional to the actual number of constellation points $N$.
-With the modulation order set as $M=32$, the optimal constellation size $N$ and the design parameter $\rho$ as a function of channel SNR $\gamma$ are illustrated in Fig.
+The proposed NENU constellation also offers a more efficient demapping implementation due to its reduced constellation cardinality.
+In the LLR computation in (15), since Euclidean distances are compared among $|\mathcal{A}(M,N,\rho)|=N$ candidate constellation points, the number of required real multiplications is given by $N$.
+With the modulation order set to $M=32$, the optimal constellation size $N$ and the design parameter $\rho$ as a function of channel SNR $\gamma$ are illustrated in Fig.
 6.
-It is observed that the required number of distinct signal points decreases as the SNR decreases.
+It is observed that the required number of distinct signal points reduces as the SNR decreases.
 It thus demonstrates that the proposed NENU constellation not only achieves near-optimal gain but also minimizes the demapping complexity.
+We note that several simplified demapping algorithms have been developed for both one-dimensional and two-dimensional NUCs [fuentes2015low, hong2020backward, hong2022enhanced, barrueco2017condensation, duan2021non, barrueco2018low].
+Quadrant search reduction (QSR) [fuentes2015low] reduces the number of the required Euclidean distance computations by discarding many symbols in different quadrants that have negligible impact on the LLR.
+Virtual point searching (VPS) [hong2020backward, hong2022enhanced], on the other hand, approximates a general two-dimensional constellation by using virtual points located on several radial axes, similar to the amplitude-phase shift keying (APSK) constellation.
+While this approximation can significantly reduce the demapping complexity, it could lead to a substantial degradation in the achievable information rate.
+Condensed symbol reduction (CSR) [barrueco2017condensation, duan2021non, barrueco2018low] reduces the demapping complexity by exploiting the clustering of constellation points that typically appears in NUCs optimized for low code rates, where each cluster is treated as a single symbol in the LLR computation.
+CSR can be applied to both one-dimensional and two-dimensional NUCs.
+In the case of two-dimensional NUCs, it is sometimes combined with VPS or QSR to further reduce the complexity [fuentes2015low, hong2020backward, hong2022enhanced].
+Nevertheless, the most effective reduction in demapping complexity is still achieved by exploiting I/Q independence, which naturally favors one-dimensional constellations especially for high modulation orders.
+In this context, a one-dimensional NUC with a carefully designed CSR-based demapper may achieve performance comparable to the proposed NENU constellation.
+However, it should rely on heuristic procedures that determine the initial constellation points and perform the clustering process.
+In contrast, since the proposed NENU constellation is uniquely described by a simple mathematical expression of (5), it ensures systematic reproducibility and enables a significantly simpler optimization process.
 IV Application to BICM-OFDM With CAF and CNC
-The design of wireless signals with low PAPR continues to be a prominent and challenging field of research [41, 42, 43, 44], and a major practical drawback of constellation shaping is its increasing PAPR as it targets Gaussian distribution.
-On the other hand, BICM-OFDM is a de facto standard for high data rate wireless communications over frequency-selective fading channels, but due to its multicarrier nature, its signal has high PAPR regardless of the constellations employed.
-Thus, PAPR reduction should be necessary for better amplifier efficiency.
-Among many PAPR reduction approaches developed for OFDM, we suggest employing CAF [13], as it is directly applicable to the standard OFDM systems.
+The design of wireless signals with low PAPR continues to be a prominent and challenging field of research [9725261, 9747919, 9962325, 10530524].
+BICM-OFDM is a de facto standard for high data rate wireless communications over frequency-selective fading channels, but due to its multicarrier nature, its signal has high PAPR regardless of the constellations employed.
+Thus, PAPR reduction should be necessary for better power amplifier efficiency.
+Among many PAPR reduction approaches developed for OFDM, we suggest employing CAF [ochiai2002performance], as it is directly applicable to the BICM-OFDM system using the proposed NENU constellation.
 The major drawback is its nonlinear distortion, which should be compensated for at the receiver.
-Therefore, we also suggest using CNC [14], which allows us to take advantage of one-dimensional constellations for efficient implementation [15].
+Therefore, we address the use of CNC [chen2003iterative, sun2021], which is also compatible with our constellation shaping.
 We demonstrate that the proposed NENU constellations can be flexibly optimized for such systems, achieving notable gains under practical scenarios.
 In what follows, for simplicity of notation, we define the set of QAM symbols composed of the two NENU PAM constellations as
-${\cal Z}_{M,N,\rho}\triangleq\{X_{I}+jX_{Q}\,|\,X_{I},X_{Q}\in{\cal A}\left(M,N,\rho\right)\}.$
+${\cal X}_{M,N,\rho}\triangleq\{A_{I}+jA_{Q}\,|\,A_{I},A_{Q}\in{\cal A}\left(M,N,\rho\right)\}.$
 IV-A OFDM Transmitter With CAF
@@
 CAF is implemented through a sequence of operations: oversampling using a $JN_{c}$-point IDFT, soft envelope limiter, and out-of-band signal removal.
-We follow the description of [13] for CAF and that of [15] for CNC with an extension to coded OFDM.
+We follow the description of [ochiai2002performance] for CAF and that of [sun2021] for CNC with an extension to coded OFDM.
 The block diagram of the OFDM transmitter with CAF is illustrated in Fig.
@@
 The codeword is bit-wise interleaved and then mapped onto the corresponding NENU constellations.
-Let $\mathbf{Z}=\left(Z_{0},Z_{1},\ldots,Z_{N_{c}-1}\right)\in\mathcal{Z}_{M,N,\rho}^{N_{c}}$ denote a vector of QAM symbols carried by each $N_{c}$-subcarrier OFDM signal, where $Z_{k}=X_{I,k}+jX_{Q,k}$ with $X_{I,k}$ and $X_{Q,k}$ representing NENU PAM symbol on the $k$th subcarrier.
-The $J$-times oversampled OFDM symbol $\mathbf{S}=(S_{0},S_{1},...,S_{JN_{c}-1})\in{\mathbb{C}}^{JN_{c}}$ is generated by performing $JN_{c}$-point inverse discrete Fourier transform (IDFT) on the zero-padded symbol vector $\mathbf{Z}^{\prime}=(\mathbf{Z},\underbrace{0,0,\ldots,0}_{(J-1)N_{c}})\in{\mathbb{C}}^{JN_{c}}$ as follows:444 In practice, DFT and IDFT are implemented using the fast Fourier transform (FFT) algorithm by setting the number of points $JN_{c}$ as the powers of two.
-$S_{n}=\frac{1}{\sqrt{JN_{c}}}\sum_{k=0}^{JN_{c}-1}Z^{\prime}_{k}e^{j2\pi k\frac{n}{JN_{c}}},\quad n=0,1,\ldots,JN_{c}-1.$ (14)
-To restrict the peak power, these time domain signal samples are clipped by soft-envelope limiter [13] as
-$\tilde{S}_{n}=\begin{cases}S_{n},&\text{ for }\left|S_{n}\right|\leq A_{\text{max}},\\ A_{\text{max}}e^{j\arg S_{n}},&\text{ for }\left|S_{n}\right|>A_{\text{max}},\end{cases}$ (15)
-where $A_{\text{max}}$ represents the maximum envelope level. We define the clipping ratio $\gamma_{\text{CR}}$ as
-$\gamma_{\text{CR}}\triangleq\frac{A_{\text{max}}}{\sqrt{P_{\text{in}}}},$ (16)
-where $P_{\mathrm{in}}\triangleq\mathbb{E}\left[\|\mathbf{Z}\|^{2}\right]/N_{c}$ is the average power of the OFDM signal prior to CAF.
+Let $\mathbf{X}=\left(X_{0},X_{1},\ldots,X_{N_{c}-1}\right)\in\mathcal{X}_{M,N,\rho}^{N_{c}}$ denote a vector of QAM symbols carried by each $N_{c}$-subcarrier OFDM signal, where $X_{k}=A_{I,k}+jA_{Q,k}$ with $A_{I,k}$ and $A_{Q,k}$ representing NENU PAM symbol on the $k$th subcarrier.
+The $J$-times oversampled OFDM symbol $\mathbf{S}=(S_{0},S_{1},...,S_{JN_{c}-1})\in{\mathbb{C}}^{JN_{c}}$ is generated by performing $JN_{c}$-point inverse discrete Fourier transform (IDFT) on the zero-padded symbol vector $\mathbf{X}^{\prime}=(\mathbf{X},\underbrace{0,0,\ldots,0}_{(J-1)N_{c}})\in{\mathbb{C}}^{JN_{c}}$ as follows: [^4] In practice, DFT and IDFT are implemented using the fast Fourier transform (FFT) algorithm by setting the number of points $JN_{c}$ as the powers of two.
+$S_{n}=\frac{1}{\sqrt{JN_{c}}}\sum_{k=0}^{JN_{c}-1}X^{\prime}_{k}e^{j2\pi k\frac{n}{JN_{c}}},\quad n=0,1,\ldots,JN_{c}-1.$ (16)
+To restrict the peak power, these time domain signal samples are clipped by soft-envelope limiter [ochiai2002performance] as
+$\tilde{S}_{n}=\begin{cases}S_{n},&\text{ for }\left|S_{n}\right|\leq r_{\text{max}},\\ r_{\text{max}}e^{j\arg S_{n}},&\text{ for }\left|S_{n}\right|>r_{\text{max}},\end{cases}$ (17)
+where $r_{\text{max}}$ represents the maximum envelope level. We define the clipping ratio $\gamma_{\text{CR}}$ as
+$\gamma_{\text{CR}}\triangleq\frac{r_{\text{max}}}{\sqrt{P_{\text{in}}}},$ (18)
+where $P_{\mathrm{in}}\triangleq\mathbb{E}\left[\|\mathbf{X}\|^{2}\right]/N_{c}$ is the average power of the OFDM signal prior to CAF.
 We then perform $JN_{c}$-point discrete Fourier transform (DFT) on the clipped signals to transform them back to the frequency-domain symbols represented by $\mathbf{U}^{\prime}=\left(U^{\prime}_{0},U^{\prime}_{1},\ldots,U^{\prime}_{JN_{c}-1}\right)\in{\mathbb{C}}^{JN_{c}}$, where
@@
 Each iteration involves estimating the transmitted symbols and applying the CAF process, including $JN_{c}$-point IDFT and DFT.
-At the receiver, we consider performing CNC to mitigate the in-band distortion caused by CAF prior to channel decoding.
+At the receiver, we consider performing CNC to mitigate the in-band distortion caused by CAF prior to channel decoding [^5] It is possible to design CNC that takes into account the channel decoder output as well as re-encoder in its loop, but this approach makes not only the receiver complexity and latency prohibitive, but also the appropriate system design more challenging.
+.
 The block diagram of the OFDM receiver with CNC is illustrated in Fig.
@@
 In the case of an AWGN channel, the received symbol corresponding to the $k$th subcarrier can be expressed as
-$\tilde{Z}_{k}=\alpha Z_{k}+D_{k}+W_{k},$ (19)
+$\tilde{Y}_{k}=\alpha X_{k}+D_{k}+W_{k},$ (21)
 where $D_{k}\in{\mathbb{C}}$ denotes the distortion component caused by the CAF process and $W_{k}\in{\mathbb{C}}$ represents the Gaussian noise with $W_{k}\sim\mathcal{CN}\left(0,N_{0}\right)$, both corresponding to the $k$th subcarrier.
-Assuming that the baseband OFDM signal follows a complex Gaussian distribution, the attenuation factor $\alpha$ due to clipping can be expressed as [13]
+Assuming that the baseband OFDM signal follows a complex Gaussian distribution, the attenuation factor $\alpha$ due to clipping can be expressed as [ochiai2002performance]
 $\alpha=1-e^{-\gamma_{\text{CR}}^{2}}+\frac{\sqrt{\pi}\gamma_{\text{CR}}}{2}\operatorname{erfc}(\gamma_{\text{CR}}),$ (22)
@@
 Provided that the length of cyclic prefix is sufficient compared to that of the channel impulse response, the received symbol over a frequency-selective fading channel can be expressed as
-$Y_{k}=H_{k}\left(\alpha Z_{k}+D_{k}\right)+W_{k},$ (21)
+$Y_{k}=H_{k}\left(\alpha X_{k}+D_{k}\right)+W_{k},$ (23)
 where $H_{k}$ is the channel coefficient of the $k$th subcarrier. Assuming that the channel coefficients can be perfectly estimated at the receiver, we may apply the zero-forcing (ZF) equalization as
-$\tilde{Y}_{k}\triangleq\frac{Y_{k}}{H_{k}}=\alpha Z_{k}+D_{k}+W_{k}^{\prime},$ (22)
+$\tilde{Y}_{k}\triangleq\frac{Y_{k}}{H_{k}}=\alpha X_{k}+D_{k}+W_{k}^{\prime},$ (24)
 where $W_{k}^{\prime}\triangleq W_{k}/H_{k}\sim\mathcal{CN}\left(0,N_{0}/\left|H_{k}\right|^{2}\right)$.
@@
   1. 1.
-    Apply the hard decision based on the constellation set $\mathcal{A}({M,N,\rho})$ to the equalized and normalized received signal $Y_{k}^{(K)}/\alpha$ (with $Y_{k}^{(0)}:=\tilde{Y}_{k}$) to estimate the transmitted symbol $\hat{Z}^{(K)}_{k}={\hat{X}_{I,k}^{(K)}}+j{\hat{X}_{Q,k}^{(K)}}$, which is given by
-    $\begin{cases}{\hat{X}}_{I,k}^{(K)}=\arg\min_{{X}\in\mathcal{A}({M,N,\rho})}\left({X}-\Re\left\{{{Y}}_{k}^{(K)}\right\}/\alpha\right)^{2},\\ {\hat{X}}_{Q,k}^{(K)}=\arg\min_{{X}\in\mathcal{A}({M,N,\rho})}\left({X}-\Im\left\{{{Y}}_{k}^{(K)}\right\}/\alpha\right)^{2}.\end{cases}$ (23)
+    Apply the hard decision based on the constellation set $\mathcal{A}({M,N,\rho})$ to the equalized and normalized received signal $Y_{k}^{(K)}/\alpha$ (with $Y_{k}^{(0)}:=\tilde{Y}_{k}$) to estimate the transmitted symbol $\hat{X}^{(K)}_{k}={\hat{A}_{I,k}^{(K)}}+j{\hat{A}_{Q,k}^{(K)}}$, which is given by
+    $\begin{cases}{\hat{A}}_{I,k}^{(K)}=\arg\min_{{A}\in\mathcal{A}({M,N,\rho})}\left({A}-\Re\left\{{{Y}}_{k}^{(K)}\right\}/\alpha\right)^{2},\\ {\hat{A}}_{Q,k}^{(K)}=\arg\min_{{A}\in\mathcal{A}({M,N,\rho})}\left({A}-\Im\left\{{{Y}}_{k}^{(K)}\right\}/\alpha\right)^{2}.\end{cases}$ (25)
   2. 2.
-    Apply the same CAF process at the transmitter to obtain the estimated in-band component, which can be modeled as
-    $\hat{{U}}_{k}^{(K)}=\alpha\hat{{Z}}_{k}^{(K)}+\hat{{D}}_{k}^{(K)}.$ (24)
+    Apply the same CAF process at the transmitter to obtain the estimated in-band component, which is modeled as
+    $\hat{{U}}_{k}^{(K)}=\alpha\hat{{X}}_{k}^{(K)}+\hat{{D}}_{k}^{(K)}.$ (26)
   3. 3.
     Calculate the distortion component by
-    $\hat{{D}}_{k}^{(K)}=\hat{{U}}_{k}^{(K)}-\alpha\hat{{Z}}_{k}^{(K)},$ (25)
+    $\hat{{D}}_{k}^{(K)}=\hat{{U}}_{k}^{(K)}-\alpha\hat{{X}}_{k}^{(K)},$ (27)
     and subtract it from the original received symbol as
     ${{Y}}_{k}^{(K+1)}$ $=\tilde{Y}_{k}-\hat{{D}}_{k}^{(K)}$
-    $=\alpha{Z}_{k}+\left({D}_{k}-\hat{{D}}_{k}^{(K)}\right)+{W}_{k}.$ (26)
+    $=\alpha{X}_{k}+\left({D}_{k}-\hat{{D}}_{k}^{(K)}\right)+{W}_{k}.$ (28)
   4. 4.
@@
 Although a lower clipping ratio helps reduce PAPR and improve power amplifier efficiency, it introduces more severe nonlinear distortion, which may prevent CNC from functioning properly.
-Note that since our NENU constellations are designed based on a one-dimensional constellation, the complexity associated with the hard decision in (23) is significantly lower than that of the two-dimensional constellation, similar to the argument of demapper complexity.
-This observation also justifies the use of one-dimensional constellations for our BICM-OFDM systems involving CAF and CNC [^5] It is possible to design CNC that takes into account the channel decoder output as well as re-encoder in its loop, but this approach makes not only the receiver complexity and latency prohibitive, but also the appropriate system design more challenging.
+Note that since our NENU constellations are designed based on a one-dimensional constellation, the complexity associated with the hard decision in (25) is considerably lower than that of a two-dimensional constellation, similar to the argument of demapper complexity discussed in Section III-E.
+In fact, if the modulation scheme is based on a general two-dimensional constellation denoted by $\mathcal{X}$, the hard decision must be performed in the complex plane as
+${\hat{X}}_{k}^{(K)}=\arg\min_{{X}\in\mathcal{X}}\left|{X}-{{Y}}_{k}^{(K)}/\alpha\right|^{2},$ (29)
+which requires comparing Euclidean distances among $\left|{\cal X}\right|=M^{2}$ symbols, instead of $\left|\mathcal{A}({M,N,\rho})\right|=N$ symbols.
+This observation also justifies the use of one-dimensional constellations for our BICM-OFDM systems employing CAF and CNC.
+IV-C Parameter Optimization for OFDM with CAF and CNC
+The performance of OFDM with CAF and CNC significantly depends on system parameters, including the number of subcarriers, clipping ratio, the number of iterations, and the channel models [sun2021].
+Therefore, it can only be evaluated by simulations for a given system setup.
+In what follows, we will first define the realistic system setup and then attempt to optimize shaping parameters based on computer simulations.
+IV-C1 System Setup
+The specific system parameters as well as the channel model considered through the remainder of this section are listed in Table III: We consider a specific OFDM system with the number of subcarriers set to $N_{c}=1024$, with each subcarrier modulated by $32$-PAM per dimension, corresponding to $1024$-QAM per subcarrier.
+Channel coding is performed using an LDPC code compliant with the 5G New Radio (NR) standard, with a codeword length of $10240$ bits [etsi38138].
+The code rate is fixed to $R_{c}=4/5$, which results in a spectral efficiency of $4.0$ bits per dimension, or $8.0$ bits per complex channel use.
+At the receiver, decoding is carried out using an offset min-sum decoding algorithm.
+Both AWGN and fading channels are considered.
+For the fading scenario, we adopt a frequency-selective Rayleigh fading channel following the 5G NR Tapped Delay Line (TDL)-B model [etsi2020138].
+In the case of CAF and CNC operations, the oversampling factor is fixed to $J=4$ throughout this work.
+Based on the preliminary simulation with the above-mentioned system setup, we found that the clipping ratio of $\gamma_{\text{CR}}=1.5$ is the minimum acceptable value that does not cause the CNC operation to malfunction [^6] Optimal clipping ratio may be mathematically analyzed in terms of signal-to-noise plus distortion power ratio (SNDR) when CAF is applied [azolini], but since our system makes use of CNC at the receiver, its error rate performance depends on several parameters such as the number of subcarriers and constellations [sun2021].
+As a result, it may become mathematically intractable.
 .
-IV-C Parameter Optimization for OFDM with CAF and CNC
-IV-C1 System Setup
-In the following numerical results, we consider the specific system where the number of subcarriers is $N_{c}=1024$ and each subcarrier is modulated by $32$-PAM per dimension ($1024$-QAM per subcarrier symbol) with $R_{c}=5/6$.
-The resulting spectral efficiency is $4.166$ bits per dimension, or $8.333$ bits per complex channel use.
-Based on the preliminary analysis, the clipping ratio of $\gamma_{\text{CR}}=1.5$ is found to be the minimum acceptable value that does not impose malfunctioning of CNC operation666 Optimal clipping ratio may be mathematically analyzed in terms of signal-to-noise plus distortion power ratio (SNDR) when CAF is applied [45], but since our system makes use of CNC at the receiver, its error rate performance depends on several parameters such as the number of subcarriers and constellations [15].
-Therefore, it may be mathematically intractable.
-.
-Although some residual clipping noise still remains as a result of CNC, the proposed system can be designed to operate properly with $\gamma_{\text{CR}}=1.5$.
-Both AWGN and fading channels are considered, where in the case of fading channels we adopt frequency-selective Rayleigh fading channel with an uncorrelated four-path equal-power delay profile as a specific example.
-The proposed NENU constellation used in each system is optimized based on the corresponding channel condition.
+TABLE III: System parameters and channel model.
+The number of subcarriers $N_{c}=1024$
+Modulation $1024$-QAM ($32$-PAM each)
+Clipping ratio $\gamma_{\text{CR}}=1.5$
+Oversampling factor $J=4$
+Maximum CNC iterations $K_{\text{max}}=10$
+Channel code 5G NR LDPC [etsi38138]
+Decoder Offset min-sum decoding
+Code rate $4/5$
+Codeword length 10240
+Spectral efficiency 4.0 per dimension
+Fading channel 5G NR Tapped Delay Line (TDL)-B [etsi2020138]
+[…]
 IV-C2 Parameter Optimization
-We optimize the system based on (10), where the BMI is computed using (8) by observing the received symbol that serves as an input to the channel decoder.
-Note that for the systems employing CAF and CNC, the channel becomes a mixture of AWGN and residual clipping noise, which may not be described by a tractable mathematical model.
-Therefore, we compute the resulting BMI based on simulations, making the optimization process more computationally intensive compared to the AWGN channels without CAF.
-We consider the following three cases:
-  1. A.
-    Without CAF, both AWGN and fading channels
-  2. B.
-    With CAF and CNC, AWGN channel
-  3. C.
-    With CAF and CNC, fading channel
-For each case, the following optimal parameter sets are found through preliminary simulations:
-  1. A.
-    $N=32$, $\rho=0.88$  (SNR = $26$ dB)
-  2. B.
-    $N=24$, $\rho=0.93$  (SNR = $27$ dB)
-  3. C.
-    $N=28$, $\rho=0.97$  (SNR = $33$ dB)
-In what follows, we refer to the NENU constellations designed by the above parameter sets as Modes A, B, and C, respectively.
-From the above results, we observe that the optimal number of constellation points depends on the channel conditions, which agrees with the trend discussed in Section III.
-Without CAF, $N=32$ is selected as optimal due to the high spectral efficiency assumption with $R_{c}=5/6$.
-In contrast, with CAF and CNC, smaller values of $N$ are preferred to better accommodate the residual nonlinear distortion that reduces the effective SNR.
-We note that constellation design based on exhaustive numerical optimization algorithms, such as particle swarm optimization (PSO), does not rely on a predefined set of constellation candidates.
-Consequently, optimizing the objective functions such as the BMI after CAF and CNC processing becomes computationally demanding.
-On the other hand, the proposed design requires evaluation over a limited number of constellation candidates determined only by the two parameters.
-It thus reduces the computational burden significantly and enables optimization in an efficient manner, even under such complicated system constraints.
-IV-C3 Total BMI
-Figure 9: Comparison of BMI values for 32-PAM constellations over an AWGN channel.
-The following three systems are evaluated: A1) Without CAF, A2) with CAF only, and A3) with CAF/CNC.
-(Parameters: $\gamma_{\text{CR}}=1.5$ for CAF and $K_{\text{max}}=10$ for CNC.)
-Figure 10: Comparison of BMI values for 32-PAM constellations over a frequency-selective Rayleigh fading channel.
-The following two systems are evaluated: F1) Without CAF and F2) with CAF and CNC.
-(Parameters: $\gamma_{\text{CR}}=1.5$ for CAF and $K_{\text{max}}=10$ for CNC.)
-We evaluate the total BMI over an AWGN channel for the following three systems:
+TABLE IV: System description and optimal parameters.
+Mode CAF/CNC Target Channel Parameters $(N,\rho)$
+A not applied AWGN/fading $(32,0.88)$
+B applied AWGN $(24,0.96)$
+C applied fading $(26,0.90)$
+[…]
+We optimize the NENU constellation for specific system configurations and channel conditions. We consider the following three system scenarios:
+  • A.
+    Without CAF, the system operates over either an AWGN channel or a fading channel.
+  • B.
+    With CAF and CNC, the system operates over an AWGN channel.
+  • C.
+    With CAF and CNC, the system operates over a fading channel.
+In systems that utilize CAF with $\gamma_{\text{CR}}=1.5$, CNC is unable to completely eliminate clipping noise.
+Consequently, the effective channel becomes a mixture of AWGN and residual clipping noise, which may not possess a mathematically tractable representation.
+In subsequent studies focused on BMI evaluation and LLR computation, the noise variance $N_{0}$ specifically represents only the AWGN components.
+This limitation occurs because the residual clipping noise is affected by the constellation geometry and the transmitted symbol sequence, making it difficult to consistently incorporate it into $N_{0}$ across different modulation schemes.
+Both the number of constellation points and their distribution should be designed to match the corresponding system condition.
+To this end, we evaluate the BER over a family of NENU constellations parameterized by $(N,\rho)$ and select the constellation that achieves a BER level of $10^{-4}$ at the lowest SNR as the representative design [^7] As discussed in Section III, the two parameters can also be selected based on the BMI.
+However, when the codeword length is not sufficient, and thus the gap between the channel capacity and the practical decoding waterfall region becomes non-negligible, BER-based optimization provides a more accurate criterion for parameter selection..
+For each system scenario, the optimal parameter set $(N,\rho)$ obtained through preliminary simulations, as well as their descriptions, are listed in Table IV.
+In what follows, we refer to the resulting NENU constellations corresponding to the above systems A, B, and C as Modes A, B, and C, respectively, and denote them by $\mathcal{X}_{\sf A}$}, $\mathcal{X}_{\sf B}$}, and $\mathcal{X}_{\sf C}$}.
+Here, we note that in the absence of CAF, the optimal parameter set $(N,\rho)$ turned out to be identical for both AWGN and fading channels, and thus they are unified as Mode A.
+From the above results, we observe that the optimal number of constellation points depends on the channel conditions, which agrees with the observation in Section III.
+Without CAF, $N=M=32$ is selected as optimal due to the high spectral efficiency setup with $R_{c}=4/5$.
+In contrast, with CAF and CNC, smaller values of $N$ are preferred to better accommodate the residual nonlinear distortion that reduces the effective SNR, but the optimal values may differ depending on the channel models.
+It is worth mentioning that constellation design based on exhaustive numerical optimization algorithms, such as particle swarm optimization (PSO), does not rely on a predefined set of constellation candidates.
+Therefore, the optimized constellation, if found, should achieve near-optimal performance.
+However, if the objective function for optimization itself is computationally demanding, e.g., the BER after CAF and CNC processing, the exhaustive search should become challenging.
+On the other hand, due to limited degrees of freedom, the proposed shaping scheme requires evaluation over a tractable number of constellation candidates.
+It thus significantly reduces the computational burden and enables efficient optimization, even if the objective function is mathematically intractable and thus requires exhaustive simulations.
+IV-C3 Total BMI for AWGN Channel
+Figure 9: Comparison of BMI values for 32-PAM constellations over an AWGN channel. The following three systems are evaluated: A1) Without CAF, A2) with CAF only, and A3) with CAF/CNC.
+Figure 10: Comparison of BMI values for 32-PAM constellations over a frequency-selective Rayleigh fading channel. The following two systems are evaluated: F1) Without CAF and F2) with CAF and CNC.
+For the purpose of investigating the effect of the parameter optimizations, we evaluate the total BMI over an AWGN channel, focusing on the following three systems:
   • A1:
-    BICM-OFDM without CAF (no PAPR reduction)
+    BICM-OFDM without CAF (no PAPR reduction).
   • A2:
-    BICM-OFDM with CAF but without CNC
+    BICM-OFDM with CAF but without CNC.
   • A3:
@@
 The results are compared in Fig.
-9, where the parameters of the NENU constellations are fixed as either Mode A or B over the entire channel SNR.
+9, where the NENU constellations are fixed as either $\mathcal{X}_{\sf A}$ or $\mathcal{X}_{\sf B}$ over the entire channel SNR.
 The uniform PAM is also shown as a reference.
-In A1, Mode A demonstrates superior performance compared to the others at the target information rate.
+In A1, the constellation $\mathcal{X}_{\sf A}$ demonstrates superior performance compared to the others at the target information rate.
 On the other hand, in A2, the achievable rates of all constellations significantly deviate from the target as a result of severe clipping noise.
 Finally, in A3, the information rate is successfully recovered by CNC, allowing the target information rate to be achieved with a lower PAPR compared to A1.
-Mode B demonstrates superior performance, yielding a greater gain over Mode A.
+The constellation $\mathcal{X}_{\sf B}$ exhibits superior performance, yielding a greater gain over the constellation $\mathcal{X}_{\sf A}$.
 The results thus demonstrate the effectiveness of the proposed constellation design tailored to the BICM-OFDM system with CAF and CNC.
-We next consider the frequency-selective Rayleigh fading channel for the following two systems:
+IV-C4 Total BMI for Fading Channel
+We next consider the fading channel (5G NR TDL-B) for the following two systems:
   • F1:
-    BICM-OFDM without CAF (no PAPR reduction)
+    BICM-OFDM without CAF (no PAPR reduction).
   • F2:
@@
 The BMI values are compared in Fig.
-10, where the parameters of the NENU constellations are fixed as either Mode A or C over the entire channel SNR, along with the uniform PAM as a reference.
-It is observed that for F1, the shaping gains achieved by the proposed constellations decrease compared to the results over the AWGN channel.
+10, where the parameters of the NENU constellations are fixed as either $\mathcal{X}_{\sf A}$ or $\mathcal{X}_{\sf C}$ over the entire channel SNR, along with the uniform PAM as a reference.
+It is observed that for F1, the gains achieved by the proposed constellations slightly decrease compared to the results over the AWGN channel.
 This may be due to the fact that a single modulation scheme is applied across all subcarriers, while they experience different effective SNRs.
 Nevertheless, in F2, significant gains by the proposed shaping are still observed when compared at the target information rate.
-Additionally, Mode C (optimized for this system over fading channel) offers enhanced performance over Mode A.
+Additionally, the constellation $\mathcal{X}_{\sf C}$ (optimized for this system over a fading channel) offers enhanced performance over $\mathcal{X}_{\sf A}$.
 V Simulation Results
-We evaluate the system-level performance of the BICM-OFDM systems employing the proposed NENU constellations in the presence of CAF and CNC.
-As performance measures, we consider the distribution of PAPR as well as BER.
-For the binary channel code, we utilize the original rate-$1/3$ turbo code with rate-compatible puncturing [46].
-The simulation setup is summarized in TABLE II.
-We compare the three systems:
+In the remainder of this paper, we present the system-level performance of the BICM-OFDM employing the proposed NENU constellations in the presence of CAF and CNC and make comparisons with other systems of similar transmitter complexity.
+The system parameters are chosen identically to Table III.
+As performance measures, we evaluate the distribution of PAPR as well as BER.
+We thus compare the following three systems:
   • S1:
@@
     BICM-OFDM employing DFT precoding.
-Note that S3 is equivalent to a single-carrier system with rectangular pulse shaping filter and thus its PAPR has direct dependence on the constellation [47].
-For each system, we evaluate the uniform PAM, the proposed NENU constellations (parameter sets: Modes A, B, and C), and the one-dimensional NUC of ATSC 3.0.
-For S3, the conventional minimum mean square error (MMSE) detector is implemented at the receiver for the frequency-selective fading channel.
-TABLE II: Simulation setup.
-OFDM 1024 subcarriers
-Modulation 32-PAM per dimension
-Coding Punctured turbo code
-Decoding Log-MAP (iteration: 15)
-Code rate 5/6
-Codeword length 30720
-Spectral efficiency $8.333$ bits per subcarrier
-PAPR reduction CAF ($\gamma_{\text{CR}}=1.5$, $J=4$) or DFT precoding
-PAPR evaluation Four-times oversampling
-CNC $K_{\text{max}}=10$
-Channel AWGN or Rayleigh fading
-Fading channel 4-path equal-power delay profile
+Note that S3 is another low-complexity PAPR reduction scheme comparable to CAF, and it is equivalent to a single-carrier system with a rectangular pulse shaping filter [myung2006single].
+Therefore, its PAPR directly depends on the underlying constellations [ochiai_wcl2012].
+For S3, a conventional minimum mean square error (MMSE) detector is employed at the receiver in the frequency-selective fading channel.
+For each system, the following modulation schemes are evaluated:
+  • The standard uniform $1024$-QAM.
+  • The proposed 1024-ary NENU constellations ($\mathcal{X}_{\sf A}$, $\mathcal{X}_{\sf B}$, or $\mathcal{X}_{\sf C}$) selected according to the system configuration.
+  • The one-dimensional (1D) $1024$-NUC of ATSC 3.0.
+  • The two-dimensional (2D) $1024$-NUC from [sillekens2022high].
+The 1D-NUC and 2D-NUC are optimized according to their respective design algorithms for the same target spectral efficiency of $8.0$ bits per complex channel use and thus are suitable for the conventional OFDM system without CAF and CNC (i.e., S1 and S3).
 V-A CCDF of PAPR
-Figure 11: Comparison of the CCDF of PAPR among the three systems: conventional OFDM without CAF, OFDM with CAF, and OFDM with DFT precoding (single-carrier).
-For each system, the three PAM constellations (uniform, Mode A, and Mode B) are evaluated.
-The CCDF curves of the PAPR among the three systems are compared in Fig. 11, where the three PAM constellations (uniform as well as NENU with Modes A and B) are compared.
-We observe no difference among the three curves plotted for S1 and S2, suggesting that each subcarrier constellation does not affect PAPR distribution for OFDM.
+Figure 11: Comparison of the CCDF of PAPR among the three systems: conventional OFDM without CAF, OFDM with CAF, and OFDM with DFT precoding (single-carrier), each employing the four different modulation schemes.
+The CCDF curves of the PAPR among the three systems are shown in Fig.
+11, where all the aforementioned $1024$-ary constellations are compared.
+We observe no difference among the four curves plotted for S1 and S2, suggesting that each subcarrier constellation does not affect PAPR distribution for OFDM.
 On the other hand, in S3, the selection of constellations significantly affects the dynamic range of the transmit signals, indicating that the use of constellation shaping increases the PAPR as expected [8227810], even though it is still less than the conventional OFDM (S1).
-In contrast, from the results of S2, we observe that CAF effectively reduces the peak power of OFDM signals, even in the presence of constellation shaping, achieving a PAPR reduction of $5.0$ dB compared to S1 when measured at a CCDF of $10^{-4}$.
+In contrast, the results for S2 indicate that CAF effectively reduces the peak power of OFDM signals even in the presence of constellation shaping, achieving a PAPR reduction of $5.0$ dB compared to S1 when measured at a CCDF of $10^{-4}$.
 V-B BER Comparison Over AWGN Channel
-Figure 12: BER comparison over an AWGN channel for the BICM-OFDM systems without CAF (no PAPR reduction) and with CAF/CNC employing three different PAM constellations.
-The BER performance over an AWGN channel for the BICM-OFDM systems with different constellations is compared in Fig.
-12, evaluating either without CAF (no PAPR reduction) or with CAF and CNC.
-As for constellations, Modes A and B are applied to $\bf S1$} and $\bf S2$}, respectively.
-In addition, the uniform PAM and one-dimensional NUC adopted in ATSC 3.0 are applied to both systems.
-(Note that the results of S3 are omitted as they exhibit similar performance to S1 over the AWGN channel.) In the case of S1, it can be confirmed that the performance gain achieved by Mode A agrees with that estimated from the BMI values in Fig.
+Figure 12: BER comparison over an AWGN channel for the BICM-OFDM systems without CAF (no PAPR reduction) and with CAF/CNC employing the four different modulation schemes.
+Fig.
+12 compares the BER performance over an AWGN channel for the BICM-OFDM systems using different modulation schemes, considering both cases without CAF (i.e., no PAPR reduction) and with CAF and CNC.
+As for the proposed shaping scheme, the constellations $\mathcal{X}_{\sf A}$} and $\mathcal{X}_{\sf B}$} are applied to $\bf S1$} and $\bf S2$}, respectively.
+Note that the results of S3 are omitted as they exhibit similar performance to S1 over the AWGN channel.
+In the case of S1, it can be confirmed that the performance gain achieved by the constellation $\mathcal{X}_{\sf A}$} agrees with that estimated from the BMI values in Fig.
 9.
-The performance is also comparable to that achieved by the one-dimensional NUC of ATSC 3.0.
-On the other hand, in the case of S2, Mode B achieves a significant gain, yielding an improvement of approximately $2.35$ dB over the uniform PAM and $0.75$ dB over the NUC when evaluated at a BER of $10^{-5}$.
-This gain, combined with the performance recovery achieved through CNC, significantly reduces the gap from S1 (without PAPR reduction).
-It is noteworthy that the performance gap between S1 with the uniform PAM and S2 with Mode B is only $0.50$ dB, while they have a significant difference in terms of PAPR.
-This suggests that the proposed NENU constellation together with CAF and CNC can significantly reduce PAPR without noticeable SNR penalty over the conventional BICM-OFDM system.
+The resulting BER performance is also comparable to that obtained by the one-dimensional NUC adopted in ATSC 3.0.
+Although the two-dimensional NUC proposed in[sillekens2022high] achieves the largest shaping gain, it requires much more demapping complexity compared with other modulation schemes based on one-dimensional constellations, as discussed in Section III.
+On the other hand, in the case of S2, the constellation $\mathcal{X}_{\sf B}$} achieves a significant gain over the standard uniform PAM.
+Since $\mathcal{X}_{\sf B}$} is optimized in the presence of CAF/CNC, the BER exhibits a steep decay with increasing SNR, indicating strong robustness to residual clipping noise.
+As a result, $\mathcal{X}_{\sf B}$} also outperforms both the one-dimensional NUC and the two-dimensional NUC, which are optimized for conventional OFDM systems without PAPR reduction, particularly at low BER levels.
+These results confirm the effectiveness of optimizing the constellation for the specific OFDM system incorporating nonlinear CAF and CNC operations.
+The shaping gain achieved by $\mathcal{X}_{\sf B}$}, assisted by the performance recovery through CNC, significantly reduces the gap relative to S1 (without PAPR reduction).
+It is noteworthy that the performance gap between S1 with the uniform PAM and S2 with $\mathcal{X}_{\sf B}$} is only about $0.50$ dB, while they have a notable difference in terms of PAPR.
+This suggests that the proposed NENU constellation together with CAF and CNC can significantly reduce PAPR without a noticeable SNR penalty compared to the conventional BICM-OFDM system.
 V-C BER Comparison Over Fading Channel
-Figure 13: BER comparison over a frequency-selective Rayleigh fading channel for the BICM-OFDM systems without CAF (no PAPR reduction), with CAF/CNC, and with DFT precoding, each employing three different PAM constellations.
-Finally, we compare BER performance among the three systems S1, S2, and S3 over a frequency-selective Rayleigh fading channel in Fig.
+Figure 13: BER comparison over the fading channel for the BICM-OFDM systems without CAF (no PAPR reduction), with CAF/CNC, and with DFT precoding, each employing the four different modulation schemes.
+Finally, we compare BER performance among the three systems S1, S2, and S3 over the fading channel (5G NR TDL-B) in Fig.
 13.
-As for constellations, Modes A and C are applied to $\bf S1$} and $\bf S2$}, respectively.
-In addition, the uniform PAM and one-dimensional NUC adopted in ATSC 3.0 are applied to all three systems for comparison.
-The performance of S2 is found effective over fading channels, achieving a gain of approximately $2.50$ dB with the proposed NENU constellation.
-On the other hand, the performance of S3 (employing an MMSE detector) is considerably inferior to that of the other OFDM-based systems.
-It is known that single-carrier signals experience BER degradation when suboptimal detectors, such as MMSE, are employed at the receiver [49, 50].
-The results thus indicate that the performance loss associated with DFT precoding is more substantial than that caused by CAF with CNC.
-Consequently, for PAPR reduction in OFDM systems utilizing constellation shaping, CAF combined with CNC is superior to DFT precoding in terms of both the PAPR reduction effect and BER performance over fading channels.
+As for the proposed shaping scheme, the constellations $\mathcal{X}_{\sf A}$} is applied to both $\bf S1$} and $\bf S3$}, whereas $\mathcal{X}_{\sf C}$} is applied to $\bf S2$}.
+It is confirmed that the proposed NENU constellation provides a shaping gain for all systems over the practical fading channel as well.
+Moreover, the performance of S3 (employing an MMSE detector) is considerably inferior to that of the other OFDM-based systems.
+It is known that single-carrier signals experience BER degradation when suboptimal detectors, such as MMSE, are employed at the receiver [nasri2009performance, tajer2010diversity].
+The results thus indicate that the performance loss associated with DFT precoding is more pronounced than that caused by CAF with CNC.
+Consequently, for PAPR reduction in OFDM systems utilizing constellation shaping, CAF combined with CNC is more advantageous than DFT precoding from the viewpoint of PAPR as well as BER performance over fading channels.
 VI Conclusion
@@
 The simulation results have demonstrated that significant performance gain can be achieved over the conventional BICM-OFDM with DFT precoding in terms of both PAPR distribution and BER.
-In our simulations, we have restricted our attention to the design of $32$-PAM (i.e., $1024$-QAM), but its extension to even higher-order constellations targeting even higher information rates is straightforward, considering the simplicity of the optimization process based on only the two distinct parameters.
+In our simulations, we have restricted our attention to the design of $32$-PAM (i.e., $1024$-QAM), but its extension to even higher-order constellations targeting even higher information rates is straightforward, considering the simplicity of the optimization process based on only the two shaping parameters.
+As final remarks, our constellation design has been optimized for the simulated systems, and its robustness against other channels or system setups should be worth investigating.
+Furthermore, it is practically important to optimize the overall system by considering the effect of power amplifier nonlinearity in the presence of CAF/CNC, which we leave as future work.